### v1

# DIRAS: Efficient LLM-Assisted Annotation of Document Relevance in Retrieval Augmented Generation

Abstract
Retrieval Augmented Generation (RAG) is widely employed to ground responses to queries on domain-specific documents. But do RAG implementations leave out important information or excessively include irrelevant information? To allay these concerns, it is necessary to annotate domain-specific benchmarks to evaluate information retrieval (IR) performance, as relevance definitions vary across queries and domains. Furthermore, such benchmarks should be cost-efficiently annotated to avoid annotation selection bias. In this paper, we propose DIRAS (Domain-specific Information Retrieval Annotation with Scalability), a manual-annotation-free schema that fine-tunes open-sourced LLMs to annotate relevance labels with calibrated relevance probabilities. Extensive evaluation shows that DIRAS fine-tuned models achieve GPT-4-level performance on annotating and ranking unseen (query, document) pairs, and is helpful for real-world RAG development. [^1] We will open-source all our codes, LLM generations, and human annotations.
DIRAS: Efficient LLM-Assisted Annotation of Document Relevance
in Retrieval Augmented Generation
Jingwei Ni1,2thanks: Equal Contributions., Tobias Schimanski2footnotemark: , Meihong Lin4, Mrinmaya Sachan1, Elliott Ash1, Markus Leippold2,3 1ETH Zürich       2University of Zurich       3 Swiss Finance Institute (SFI) 4 University of Electronic Science and Technology of China {jingni, msachan, ashe}@ethz.ch, meihong_lin@uestc.edu.ch {tobias.schimanski, markus.leippold}@df.uzh.ch
1 Introduction
RAG has become a popular paradigm for NLP applications Gao et al. (2024). One core phase of RAG systems is Information Retrieval (IR), which leverages cheap retrievers to filter relevant information and thus save LLM inference costs. Given its cost-saving nature, IR might be a performance bottleneck for RAG (Chen et al., 2023b; Gao et al., 2024). Both leaving out important relevant information (low recall) as well as including excessively related but irrelevant information (low precision) may lead to severe performance down-grades (Ni et al., 2023; Cuconasu et al., 2024; Niu et al., 2024; Schimanski et al., 2024a). However, evaluation results on general-domain benchmarks (Thakur et al., 2021) may hardly indicate the IR performance on RAG systems, as the definition of relevance varies drastically across different domains and use cases (Bailey et al., 2008). See App. A for an example where the relevance judgment may vary widely with or without domain expertise. Therefore, domain-specific benchmarks need to be annotated to evaluate RAG systems in different domains.
To address this scarcity of domain-specific IR benchmarking, we propose DIRAS, a pipeline to annotate domain-specific IR data at scale (illustrated in Fig. 1). Specifically, users only need to input (1) some domain-specific queries and documents and (2) definitions for what is (ir)relevant for each query. Then, DIRAS distills relevance prediction data from GPT-4 to fine-tune open-sourced LLMs, which can then be used to annotate relevance on a large scale without extra API cost, avoiding annotation selection bias (Thakur et al., 2021) with limited expenditure.
Prior work distills open-sourced LLM-based rerankers using pairwise (Qin et al., 2024) or listwise (Sun et al., 2023b; Pradeep et al., 2023) as the ranking paradigm. However, we choose a pointwise approach (i.e., predicting a relevance score per document and then ranking according to the scores) for DIRAS, to fulfill our three Desiderata for a document relevance annotator:
D1. Efficient and Effective: When benchmarking IR of RAG systems, it is important to annotate all (query, paragraph) pairs to avoid annotation selection bias, since information thoroughness is critical for many queries (e.g., overall assessment, Ni et al., 2023). Therefore, the pointwise method is more favorable as it largely outperforms list- or pairwise method in efficiency (Sun et al., 2023a). Prior work worries that pointwise ranking is efficient but not effective due to the difficulty in calibration (Sun et al., 2023a; Qin et al., 2024). However, this method lacks empirical investigation in prior work. In our work, we observe both GPT-4 and DIRAS fine-tuned LLMs achieve very competent pointwise ranking results thanks to good calibration.
D2. Improved Leverage of the Relevance Definition: Relevance and partial relevance judgments might be subjective without domain expertise or explicit annotation guidelines (Bailey et al., 2008; Saracevic, 2008; Thomas et al., 2024; see also App. A). Therefore, DIRAS explicitly puts relevance definition into relevance prediction prompts to achieve more objective and consistent results. Compared to the list- or pairwise method, the pointwise method is provided with only one document each time. Thus, it may analyze the document along the relevance definition in better detail, especially with CoT prompting (Wei et al., 2022).
D3. Richer Predictions for RAG Requirements: Listwise or pairwise ranking algorithms only predict a relative rank for documents. With only the rank, retrieving the same number of documents (top-k) for all questions is suboptimal as it is likely that different questions have different amounts of relevant information (details in § 5). In contrast, the pointwise method predicts not only ranks but also binary relevance and calibrated relevance probabilities. Both binary labels and relevance scores allow RAG systems to retrieve the actual amount of relevant information to a question. Furthermore, the calibrated relevance probability is also helpful in automatic annotation (Ni et al., 2024), indicating which annotations are partially relevant and/or check-worthy.
To evaluate DIRAS, we conduct experiments with two datasets. First, we annotate a high-quality dataset based on ChatReport [^2] https://reports.chatclimate.ai/ (Ni et al., 2023), a real-world RAG application analyzing lengthy corporate reports. This dataset also incorporates the ideas of partial relevance and labeling uncertainty. ChatReport is representative of RAG applications that are sensitive to retrieval results, as thoroughly analyzing disclosed information is crucial for assessing what is under-addressed in the reports. Evaluation on ChatReport data shows that the pointwise ranking of DIRAS achieves very satisfactory performance, even superior to the widely-adopted listwise method (Pradeep et al., 2023) (§ 3.1). The best DIRAS fine-tuned model also achieves GPT-4-level performance in terms of relevance ranking and calibration (§ 3.2).
Second, we use a DIRAS fine-tuned model to re-annotate all (query, document) combinations (43k in total) in ClimRetrieve (Schimanski et al., 2024b), a real-world record of experts’ information seeking workflow. Experiments show that DIRAS fine-tuned models successfully understand fine-grained degree of relevance (§ 4.1) and enhance their performance through improved relevance definitions (§ 4.2). Furthermore, it mitigates IR annotation bias by identifying information ignored by experts (§ 4.3), and benchmarks IR algorithms’ target domain performance upon all 43k (query, document) pairs (§ 4.4). Finally, we propose recommendations for future RAG designs based on our takeaways (§ 5).
Collectively, our contributions include:
  1. 1.
    We propose DIRAS, a framework tuning open-sourced LLMs into efficient and effective IR annotators, taking domain expertise into account.
  2. 2.
    We annotate a high-quality IR benchmark based on ChatReport with (partial) relevance labels and uncertainty labels, based on explicit relevance definitions.
  3. 3.
    We compare DIRAS fine-tuned models with a real-world information-seeking workflow by experts, showing the model accurately understands granular relevance definitions and helps mitigate annotation selection bias.
Figure 1: DIRAS pipeline. RAG-specific queries, documents as inputs; calibrated small-LLM annotators as outputs.
2 DIRAS
2.1 DIRAS Pipeline
To address the outlined desiderata (D1, D2, D3) in § 1, we design the DIRAS pipeline (illustrated in Fig. 1). This pipeline comprises the training data creation and fine-tuning of small LLMs to the calibrated annotators.
Training Data Creation: We create the training data from domain-specific sources like reports. As a result, the data will be composed of a set of domain-specific (query, document) pairs. Each query comprises a question and a definition indicating what is relevant or irrelevant to the question. The relevance definition can be designed by human experts or generated by LLMs following App. B. To obtain the documents for each question/query, we employ a sampling strategy using top-k relevant documents ranked by a small dense retriever. We set a top-k threshold and sample an equal number of documents within and outside of the threshold for each question. While sampling in top-k aims at covering some relevant documents, sampling outside of top-k ensures covering the broader distribution of (query, document) pairs.
Open-Sourced LLM Fine-Tuning: Once we sampled the domain-specific (query, document) pairs, we create instruction fine-tuning (IFT) data with a SOTA generic LLM $\mathcal{M}$ and the prompt template $\mathcal{P}$ (see e.g., Fig. 2). The prompt template incorporates a (query, document) pair and an instruction to predict a binary relevance label and its confidence score. Finally, the distilled IFT data is used to fine-tune open-sourced LLMs $\mathcal{M}_{o}$ (“o” stands for “open”) to conduct binary relevance prediction with confidence.
2.2 Evaluation Metrics
Both $\mathcal{M}$’s and $\mathcal{M}_{o}$’s results should be evaluated against two types of human labels. (1) Relevance labels: whether a document is helpful for answering a query or not, disagreement between different annotators needs to be resolved to obtain the final relevance label. (2) Uncertainty labels: an annotation is uncertain, if the annotators have a strong disagreement, or the majority of them agree that the (query, document) is partially relevant. We use the following metrics for evaluation:
Binary Relevance: We compute the F1 Score of models’ binary relevance prediction using relevance labels. For RAG systems, binary relevance labels are important for deciding which documents should be passed to LLMs.
Calibration: Both $\mathcal{M}$ and $\mathcal{M}_{o}$ give confidence scores, which should calibrate the binary accuracy to indicate whether the prediction is trustworthy. We use Expected Calibration Error (ECE), Brier Score, and AUROC to measure calibration performance, following Kadavath et al. (2022) and Tian et al. (2023).
Information Retrieval: The confidence scores also give a calibrated relevance probability which can be used to rank documents for each query. To directly evaluate the ranking performance, we measure nDCG and MAP upon relevance labels.
Uncertainty: If the models understand the difficulty and uncertainty caused by partial relevance, they should have lower confidence scores on samples that humans found to be uncertain. Thus we compute average precision (AP) scores between confidence and uncertainty labels. We chose AP because it is threshold-free by summarizing F1 scores of all thresholds of relevance scores. Details of computing all metrics are in App. C.
3 Experiments on ChatReport
To the best of our knowledge, there is no existing IR dataset that (1) accompanies each query with domain-specific relevance definition; and (2) provides annotations of human uncertainty caused by partial relevance. Therefore, we annotate a dataset based on ChatReport (Ni et al., 2023) to fulfill evaluation purposes in § 2.2. ChatReport is an online RAG application for corporate climate report analyses and Q&A about the reports. [^3] See https://reports.chatclimate.ai/.
Data Sampling: We sample 31 questions and 80 reports from this application. Climate reports are sampled randomly from openly accessible user submissions [^4] See https://github.com/EdisonNi-hku/chatreport. (PDF parsing details in App. D). The questions are strategically sampled to ensure representativeness and diversity. Specifically, 11 questions are the core questions used in ChatReport, which cover essential topics of sustainability disclosure. 20 questions are selected from users’ customized questions posed to the ChatReport tool. Finally, we prompt GPT-4 to draft relevance definitions for all questions (see App. B).
Train-Test Split: We split the questions into 11 for testing and 20 for training. Similarly, we split the reports into 30 for testing and 50 for training. This ensures the evaluation on unseen questions and reports. For each question, we randomly sample 60 documents – 30 each from in the top-5 and outside the top-5 (using OpenAI text-embedding-3-small as the dense retriever). Ultimately, (query, document) pairs in training split are used to create training data with relevance label and confidence score predictions (details in § 3.1). Data points in the test split are passed to human annotation.
Test Data Annotation: We leverage relevance definitions as the annotation guidelines. A document is relevant if and only if it fulfills the definition. The data labeling process follows two steps. First, we employ two annotators who independently annotate all test data to be either relevant, irrelevant, or partially relevant. Second, we employ a subject-matter expert in corporate climate disclosure to resolve conflicts to obtain final relevance labels. Besides relevance labels, we also obtain uncertainty labels from human annotations: Whenever there is strong disagreement (co-existence of relevance and irrelevance labels) or agreement on partial relevance (two or more annotators agree on partial relevance), the data point is labeled as uncertain. Inter-annotator agreement and other details can be found in App. E.
Prompt: <question>: What is the firm’s Scope 3 emission? <question_definition>: This question is looking for information about the firm’s emission in … <paragraph>: {one text chunk from a climate report} Is <paragraph> helpful for answering <question>? Provide your best guess, and confidence score from 0 to 1. Teacher LLM $\mathcal{M}$: [Reason]: {Reason why the paragraph is (un)helpful.} [Guess]: {Yes or No.} [Confidence]: {confidence score between 0.0 and 1.0.}
Figure 2: Our prompt template $\mathcal{P}$ for distilling training data from $\mathcal{M}$. It is shortened for presentation. Full $\mathcal{P}$ is in Fig. 9.
3.1 How to Distill Training Data?
Setting Unc. Bin. Cal. Info. Avg.
List-2/1 - - - 76.86 -
List-2/1-D - - - 74.72 -
List-10/5 - - - 84.74 -
List-10/5-D - - - 84.45 -
List-20/10 - - - 78.05 -
List-20/10-D - - - 82.54 -
Point-Ask 39.27 84.07 94.41 87.57 76.33
Point-Ask-Prob-D 44.74 84.52 93.72 88.39 77.84
Point-Tok-D 28.83 86.32 93.31 80.90 72.52
Point-Ask-D 54.01 86.32 94.41 88.48 80.80
Table 1: GPT-4’s performance on ChatReport test set with different ranking methods (Point- or Listwise), with/without relevance definition (D), and calibration method (Ask or Tok). Unc. denotes the Average Precision (AP) of predicting uncertainty; Bin. denotes F1 score of binary relevance prediction; Cal. is the average of AUROC, ECE, and Brier Score; Info. is the average of nDCG and MAP; Avg. denotes the average of all metrics. Best scores of each column are bolded.
To train better open-sourced LLMs $\mathcal{M}_{o}$, it is crucial to distill high-quality training data from teacher LLM $\mathcal{M}$. Specifically, we compare the following three implementation choices:
Pointwise vs. Listwise: The listwise method is popular in ranking data distillation given its moderate cost and good performance (Sun et al., 2023b; Pradeep et al., 2023). However, the more efficient pointwise method is under-explored in prior work – majorly due to the concern about poor calibration (Sun et al., 2023a; Qin et al., 2024). Calibration method (Tok vs. Ask): One calibration method is to get the relevance confidence by probing the model’s generation probability of the token Yes/No when predicting a document’s relevance (Tok, Liang et al., 2023). An alternative way is directly asking LLMs to verbalize confidence score, which may work better for instruction following LLMs (Ask, Tian et al., 2023). With vs. without relevance definition: As ChatReport test data is annotated based on the relevance definition, performance should increase if the model correctly takes the in-context relevance definition into consideration.
Following the takeaways of Thomas et al. (2024), we design the prompt $\mathcal{P}$ for the pointwise method with detailed task/role description, relevance definition and CoT prompting (see Fig. 2 and Fig. 9, prompt without definition in Fig. 11). We use the listwise ranking prompt from Sun et al. (2023b) and Pradeep et al. (2023) (see prompt with/without definition in Fig. 13/Fig. 12). For the pointwise method, we run one variation to test prompt sensitivity: directly asking for relevance probability instead of confidence for guess (prompt in Fig. 10). As the listwise ranking is sensitive to window/step size, we run three variations with window/step sizes of 2/1, 10/5, and 20/10. text-embedding-3-small is used for listwise methods’ initial ranking. The results are shown in Table 1. We observe that: (1) With the proper calibration method (Ask), the pointwise method outperforms the listwise method. (2) The listwise method is sensitive to window size, while the pointwise method gives more consistent performance across prompts. (3) Adding relevance definition drops the listwise performance in 2 out of 3 cases, while that improves the pointwise performance. Thus we choose pointwise to be our distillation strategy.
3.2 Open-Sourced LLM Fine-Tuning
Setting Unc. Bin. Cal. Info. Avg.
Small-embed - - 83.63 66.34 -
Large-embed - - 84.07 69.36 -
BGE-Gemma - - 78.21 68.47 -
GPT-3.5 29.71 45.27 87.79 74.16 59.23
GPT-4 54.01 86.32 94.41 88.48 80.80
Llama3-CoT-Ask 36.57 76.58 93.32 86.15 73.16
Llama3-CoT-Tok 41.74 76.58 90.82 85.96 73.78
Llama3-Ask 40.18 82.11 94.04 86.02 75.59
Llama3-Tok 41.60 82.11 94.41 89.19 76.83
Phi3-CoT-Ask 36.08 72.95 92.35 80.56 70.48
Phi3-CoT-Tok 35.49 72.95 88.15 80.64 69.31
Phi3-Ask 32.30 73.23 91.16 80.05 69.18
Phi3-Tok 38.00 73.23 92.05 86.94 72.55
Gemma-CoT-Ask 31.60 72.38 91.14 81.39 69.13
Gemma-CoT-Tok 39.03 72.38 88.58 80.33 70.08
Gemma-Ask 25.74 67.13 88.27 77.43 64.64
Gemma-Tok 50.72 67.13 92.44 81.17 72.87
Table 2: Comparison between the fine-tuned $\mathcal{M}_{o}$ and different baselines on ChatReport test data. Unc. denotes the Average Precision (AP) of predicting uncertainty; Bin. denotes F1 score of binary relevance prediction; Cal. is the average of AUROC, ECE, and Brier Score; Info. is the average of nDCG and MAP; Avg. denotes the average of all metrics. The best scores are bolded and the second bests are underlined.$\dagger$ denotes the best score achieved by each LLM architecture.
DIRAS data fine-tuned models $\mathcal{M}_{o}$ will be used to predict all (query, document) combinations to mitigate annotation selection bias (Thakur et al., 2021). Therefore, shorter generations from $\mathcal{M}_{o}$ (e.g., without CoT) are favored as the inference cost for transformers increases quadratically with generation length. Additionally, the choice of calibration method matters for LLMs (Tian et al., 2023). To explore these aspects, we fine-tune $\mathcal{M}_{o}$ in four settings: $\mathcal{M}_{o}$-CoT-Ask, $\mathcal{M}_{o}$-CoT-Tok, $\mathcal{M}_{o}$-Ask, $\mathcal{M}_{o}$-Tok, where CoT means $\mathcal{M}_{o}$ is tuned to generate [Reason], [Guess], and [Confidence]; without CoT denotes $\mathcal{M}_{o}$ is tuned to only generate [Guess] and [Confidence]; “Ask” means the result is calibrated by the generated confidence score in [Confidence] field; and “Tok” means we take the token-level probability of “Yes/No” after “[Guess]:” as the confidence score for calibration. The prompt in Fig. 2 is used for fine-tuning. The “[Reason]:” line is removed in settings without CoT.
We fine-tune Llama-3-8B-instruct (AI@Meta, 2024), gemma-7b-it (Team et al., 2024b), and Phi-3-mini-4k-instruct (Abdin et al., 2024) (details in App. F). We compare these fine-tuned models with baselines including GPT-3.5 and GPT-4 using prompt $\mathcal{P}$; the OpenAI embedding models text-embedding-3-small, and text-embedding-3-large; and BGE Gemma reranker [^5] https://huggingface.co/BAAI/bge-reranker-v2-gemma, a popular LLM-based reranker for general domain. As Fig. 3 shows, fine-tuning improves original models in all settings. Furthermore, Table 2 shows the results of all fine-tuned models in comparison to all baselines. We observe that $\mathcal{M}_{o}$-Tok outperforms other settings for all LLM architectures. The best setting Llama3-Tok achieves GPT-4 level performance in calibration and IR on unseen questions and reports.
Interestingly, we find that omitting the chain of thought usually leads to a performance increase for all three LLM architectures. CoT sometimes leads to a limited increase when asking for calibration (Ask), but constantly results in a performance drop when calibrated with token-level probability (Tok). Therefore, $\mathcal{M}_{o}$ should be fine-tuned without CoT for performance and inference efficiency. Moreover, Tok rarely underperforms Ask, different from Tian et al. (2023)’s finding and our observations in Table 1. Thus, future work may consider probabilities of important tokens (e.g., Yes/No in our prompt template) as a promising calibration tool.
Figure 3: Shaded bars denote the performance of original models. Colored bars denote the improvement brought by fine-tuning.
4 Experiments on ClimRetrieve
In this section, we showcase how the DIRAS fine-tuned $\mathcal{M}_{o}$ can assist the IR annotation in a real-world setting, leveraging the ClimRetrieve dataset (Schimanski et al., 2024b). This dataset records analysts’ real-life procedure of sustainability report analyses: raise questions about sustainability reports and then go through these reports to find relevant information to answer their questions. The dataset comprises 43k (query, document) pairs, among which 595 pairs are annotated as relevant. Relevant documents are annotated with a relevance score from 1 to 3 (1 translates to partially relevant and 2/3 means relevant). App. I gives a detailed overview of the dataset.
Compared to ChatReport data, ClimRetrieve is a more challenging and realistic setting because: (1) For ChatReport data, we draft explicit relevance definitions and annotate relevance dependent on them. But for ClimRetrieve, the analysts’ mental model for what is (ir)relevant for their posed questions is unknown to us. (2) ClimRetrieve records human analysts’ real-life workflow of reading the full reports and searching for relevant information, which differs from ChatReport data where annotators are presented with (query, document) pairs as separate data points. Thus, ClimRetrieve only has gold labels for relevant (query, document) pairs. Other not annotated (query, document) combinations might be either irrelevant or a part of annotation selection bias – a widely existing problem in IR annotation (Thakur et al., 2021).
This challenging and realistic setting of ClimRetrieve allows us to investigate the following research questions regarding DIRAS fine-tuned $\mathcal{M}_{o}$: (1) RQ1: Can $\mathcal{M}_{o}$ understand fine-grained differences in degree of relevance? (2) RQ2: Can we improve $\mathcal{M}_{o}$’s performance through improving relevance definitions? (3) RQ3: Can $\mathcal{M}_{o}$ assists in mitigating annotation selection bias (Thakur et al., 2021)? (4) RQ4: Can $\mathcal{M}_{o}$’s predictions help benchmarking IR algorithms? We use the best $\mathcal{M}_{o}$ in § 3 (Llama-3 without CoT) to study these RQs.
4.1 RQ1: Understanding Fine-Grained Relevance Levels
Setting nDCG nDCG@5 nDCG@10 nDCG@15
Random 71.04 50.88 52.77 54.45
Small-embed 74.52 61.28 60.36 61.69
Large-embed 76.30 63.13 63.36 64.67
GPT-3.5 74.62 60.08 61.49 61.91
GPT-4 75.55 60.89 63.23 65.26
Llama3-Ask 77.23 67.60 66.18 67.57
Llama3-Tok 76.55 67.20 66.23 65.83
Table 3: Performance on ranking the relevant (query, document) pairs in ClimRetrieve.
We first evaluate fine-tuned Llama-3 on 595 gold labels of ClimRetrieve to verify whether it can effectively recover analysts’ ranking for relevant content by understanding which documents are more helpful than others. Relevance definitions are drafted with GPT-4 with the same procedure as § 3. We report nDCG [^6] MAP can only measure binary relevance and since we only investigate relevant samples, it cannot be calculated. scores to measure the ranking performance on ClimRetrieve. Gold labels 1, 2, and 3 are assigned with relevance scores 1/3, 2/3, and 1. Besides OpenAI 3rd generation embedding models, we also have a random baseline where all (query, document) pairs are assigned a random relevance score between 0 and 1. The random baseline results are averaged over 5 random seeds (40 to 44). Importantly, all ClimRetrieve annotations are to some degree relevant, so improvement over the random baseline is challenging as the system needs to understand the trivial different degrees of relevance.
Table 3 presents different systems’ performance. There is a clear trend of outperformance of the fine-tuned Llama-3 models in this challenging setting. They also exceed the random baseline by a significant margin, indicating the model correctly understands the fine-grained levels of relevance.
4.2 RQ2: Improving Performance through Improving Definitions
Setting nDCG MAP
[…]
Table 4: Comparison of using the generic and the expert-informed relevance definitions for ranking all ClimRetrieve (query, document) pairs.
We then test whether we can improve $\mathcal{M}_{o}$’s performance by improving the relevance definition presented in prompts. In § 3 and § 4.1, we use GPT-4 to draft the relevance definitions. To investigate the role of the definitions, we compare two setups: (1) The generic relevance definition: the definition drafted by GPT-4. (2) The informed relevance definition: we use the labeled texts in the ClimRetrieve dataset as examples to create artificial expert- textitinformed relevance definitions. This way, we simulate the inclusion of expert knowledge to improve the relevance definition (see App. J for details).
After creating the definitions, we repeat predicting the relevance scores with the DIRAS fine-tuned Llama-3. First, we repeat the setup of previous § 4.1 with only human-annotated 595 relevant (query, document) pairs. We find the utilization of expert-informed definitions produces similar results, at most improving nDCG (see App. K). We argue that the definition might only make a significant difference when predicting all (query, document) pairs instead of only the human-annotated ones. The inclusion of specific nuances might especially help when compared with non-relevant documents. Calculating the nDCG and MAP score for all 43k (query, document) pairs, we find evidence for this notion (see Table 4). Thus, the inclusion of expert-informed definitions seems to improve the performance (for more details, see App. K).
4.3 RQ3: Mitigating Annotation Bias
ClimRetrieve employs a real-world analyst scenario. This entails that the human only selectively annotates documents that are likely to be relevant and assumes unannotated documents as irrelevant (see e.g., Thakur et al., 2021). Therefore, the dataset allows us to investigate our model’s capabilities to counteract biases. For this purpose, we reannotated 288 strategically sampled (sampling details in App. M) (document, query) pairs to investigate Llama3-Tok’s performance on ClimRetrieve’s missing annotation. For samples not labeled as relevant in ClimRetrieve but by Llama3-Tok, the reannotations indicate that we can effectively overrule these blind spots. As Fig. 4 shows, increasing the threshold of Llama3-Tok confidence allows us to be increasingly sure that we indeed find a relevant document. For the around 60% of the samples with a Llama3-Tok confidence higher than 99%, almost all samples are reannotated as relevant. Besides annotation selection bias, there are also other reasons for overruling the initial annotations. ClimRetrieve annotators may rather have focused on hard relevant documents omitting partially relevant ones. Furthermore, our GPT-4 created definitions may broaden the scope of what is relevant. However, we find that our model is well-calibrated and consistent with the provided relevance definitions. Therefore, Llama3-Tok is helpful for mitigating annotation selection bias (for a detailed analysis, see App. M).
Figure 4: This table shows the percentage of agreement between Llama3-Tok and our human reannotation, and the amount of data remaining, when adjusting the confidence threshold for (query, document) pairs that were considered relevant by Llama3-Tok but not by the ClimRetrieve annotation.
4.4 RQ4: Using $\mathcal{M}_{o}$ to Benchmark IR
Setting Kendall’s $\mathbf{\tau}$
BGE-Base 35.71
BGE-Base-ft 36.34
BGE-Large 34.74
BGE-Large-ft 36.55
Table 5: Different embedding models’ performance benchmarked by $\mathcal{M}_{o}$’s prediction on all 43K (query, document) pairs of ClimRetrieve. “ft” denotes the model is fine-tuned on in-domain data.
After validating $\mathcal{M}_{o}$’s performance, we can use it to annotate all 43k ClimRetrieve datapoints and obtain a benchmarking dataset to select IR algorithms. This approach can be especially helpful when lacking human annotation and annotation selection bias is prevalent. Specifically, we compare the performance of embedding models before and after in-domain fine-tuning. If the $\mathcal{M}_{o}$-annotated benchmark gives higher scores to the fine-tuned checkpoints, that means it is capable of selecting a better model for this specific domain.
For this experiment, we first fine-tune open-sourced embedding models bge-large-en-v1.5 and bge-base-en-v1.5 (Chen et al., 2024) on ChatReport test set [^7] We fine-tune on the test instead of the training set to (1) leverage high-quality human annotation for fine-tuning; and (2) avoid indirect data leakage as $\mathcal{M}_{o}$ is fine-tuned on the training set. (fine-tuning details in App. L). We then compare embedding models’ relevance ranking with the predicted ranking of Llama3-Tok on all 43K (query, document) pairs in ClimRetrieve. We use Kendall’s $\tau$ as the metric, which directly compares the correlation between two ranks. The results are shown in Table 5. We find the Llama3-Tok-annotated benchmark successfully picks out the fine-tuned checkpoints, showing a capability of benchmarking information retrieval algorithms.
Interestingly, the unfine-tuned BGE-Base correlates more to Llama3-Tok compared to BGE-Large, although the latter shows stronger performance on MTEB (Muennighoff et al., 2023). This indicates the necessity of domain-specific benchmarking to tell the in-domain performance.
5 Recommendation for Future RAG
Avoiding Top-K Retrieval: Naive RAG systems (Ni et al., 2023) usually retrieve top-k (a fixed number k) documents to augment LLM generation. However, different questions tend to have different amounts of relevant information.Advanced RAG employs query routers to pick retrieval strategies (Gao et al., 2024). However, choosing the proper k without access to full documents is still hard. To demonstrate this, we average the relevance score (predicted by Llama3-Tok) over all documents for each question in ClimRetrieve. The resulting average relevance score will be a proxy for the amount of relevant information on the question in all reports. As Fig. 5 shows, different questions vary considerably in the amount of relevant information. Therefore, we suggest not using top-k IR, avoiding the prior determined k that does not fit the actual amount of relevant information.
Given the calibrated prediction of DIRAS fine-tuned $\mathcal{M}_{o}$, an alternative way is to retrieve all documents whose relevance scores exceed a pre-defined threshold. Thus, different questions can retrieve different amounts of information depending on whether passing the relevance threshold. Advanced RAG designs can even strategically pick the calibrated threshold for different questions, for example, allowing more partial relevance for summary queries. Fig. 6 shows the F1 Scores of GPT-4 and Llama3-Tok with different relevance thresholds. Llama3-Tok achieves good F1 scores over a wide range of thresholds. Thanks to its compact size (8B), it can be efficiently deployed as a reranker in RAG systems.
Optimizing Relevance Definitions: Results in Table 2 and Table 3 are obtained with GPT-4-drafted relevance definitions (i.e., relevance definitions). Although this approach is useful in large-scale applications, there is still space for improvement by optimizing relevance definition, as shown in § 4.2. According to Bailey et al. (2008), the question originators are the gold standard for relevance definition. Hence, with the help of DIRAS, future RAG systems may allow users to customize their requirements for relevant information.
Figure 5: The proximate amount of relevant information for 16 questions in all ClimRetrieve reports, according to Llama3-Tok’s relevance scores.
Figure 6: Instead of always retrieving top-k, we can retrieve documents if they have relevance scores higher than a threshold. This figure shows the change of F1 scores for obtaining relevant documents by thresholds.
6 Related Literature
IR plays an important role in RAG but also becomes a performance bottleneck (Gao et al., 2024). Low precision in IR may cause LLMs to hallucinate or pick up irrelevant information (Cuconasu et al., 2024; Schimanski et al., 2024a). Low recall may leave out critical information for analysis (Ni et al., 2023). Prior work proposes various datasets and algorithms to benchmark RAG performance (Saad-Falcon et al., 2023; Chen et al., 2023a; Niu et al., 2024), but most of them focus on the final generation quality. No previous work studies how to efficiently generate domain-specific IR benchmarks for RAG. Sun et al. (2023b, a); Pradeep et al. (2023); Qin et al. (2024) find that SOTA generic LLMs are good rerankers and such ability can be distilled to open-sourced LLMs. However, they all focus on pairwise or listwise ranking methods, while our work shows that the pointwise method (1) better fulfills the need of annotating domain-specific IR; and (2) works better with proper calibration method (Tian et al., 2023). Our work follows the stream of research annotating document relevance with LLMs (Thomas et al., 2024), taking one step further by (1) showing how to annotate with small LLMs; (2) predicting relevance scores to address partial relevance; and (3) use them to benefit RAG development.
7 Conclusion
In this work, we introduce the DIRAS pipeline to fine-tune open-source LLMs to calibrated annotators. We prove the effectiveness of the approach on two dataset sets. The DIRAS approach has two significant advantages: (1) it is case-specialised allowing the incorporation of domain-specific knowledge into definitions, and (2) it helps to efficiently label a huge amount of documents with calibrated relevance scores.
Limitations
As with every work, this has limitations. Our first limitation stems from the usage of two datasets focusing on a scenario of RAG report analyses. Given our expertise, this allows us to extend the deepness of our investigations: annotating domain-specific benchmarks and conducting error analyses. However, this limits the wideness of our research. While we argue that the results are representative of other knowledge-intensive RAG scenarios, it remains an open question for future work to generalize the DIRAS pipeline.
Second, this project focuses on text documents. This means we do not evaluate the performance of the DIRAS pipeline on graph and table content. While this also presents a general limitation of modern-day RAG systems, we believe it is a crucial future step to generalize DIRAS’s idea of scalable information retrieval benchmarking to multi-modality.
Our third limitation, and also a viable option to address multimodality, lies in the recent introduction of long-context LLMs. These may make the role of information retrieval in RAG less crucial as entire documents can be used to answer a question. At the same time, we observe that long-context models are good in needle-in-a-haystack problems but not as good when multiplied needles exist (Team et al., 2024a). Thus, even for long-context LLMs, an efficient system like DIRAS could enable improving algorithms for finding and using multiple relevant pieces of information or help improve the model’s ability to do so.
Ethics Statement
Human Annotation: In this work, all human annotators are Graduate, Doctorate researchers, or Professors who have good knowledge about scientific communication and entailment. They are officially hired and have full knowledge of the context and utility of the collected data. We adhered strictly to ethical guidelines, respecting the dignity, rights, safety, and well-being of all participants.
Data Privacy or Bias: There are no data privacy issues or biases against certain demographics with regard to the data collected from real-world applications and LLM generations. All artifacts we use are under a creative commons license. We also notice no ethical risks associated with this work
Reproducibility Statement: To ensure full reproducibility, we will disclose all codes and data used in this project, as well as the LLM generations, GPT-4 and human annotations. For OpenAI models, we use “gpt-4-0125-preview” and “gpt-3.5-turbo-0125”. We always fix the temperature to 0 when using APIs.
Acknowledgements
This paper has received funding from the Swiss National Science Foundation (SNSF) under the project ‘How sustainable is sustainable finance? Impact evaluation and automated greenwashing detection’ (Grant Agreement No. 100018_207800).
Appendix A Exemplifying Partial Relevance
When labeling whether a document is relevant for a question, there exists a large grey scale of relevance rather than a black-and-white relevant or irrelevant label. Humans can only consistently capture these nuances to a certain extent. The judgment of relevance also depends on the context and the annotator’s domain expertise.
Consider for instance the following excerpt of a document:
[...] Implement Risk Controls: Integrated Management System (IMS): The K&S Integrated Management System (IMS), which has been implemented at our six major design and manufacturing sites, is certified under the corporate ISO 9001:2015, ISO 14001:2015 and ISO 45001:2018 certifications. Our integrated Quality, Environmental and Occupational Health & Safety (QEHS) Management System enables the achievement of harmonized K&S worldwide objectives.
Furthermore, conclude as to whether this is relevant to answer the following question and definition:
Meaning of the question: The question "What processes does the organization use to identify and assess climate-related risks?" is asking for information about the specific methods, tools, or strategies that a company employs to recognize and evaluate the potential risks to its operations, financial performance, and overall sustainability that are associated with climate change. This includes understanding how the organization anticipates, quantifies, and plans for the impacts of climate-related phenomena such as extreme weather events, long-term shifts in climate patterns, and regulatory changes aimed at mitigating climate change.
Examples of information that the question is looking for:
1. The use of climate risk assessment tools or software that helps in modeling and predicting potential impacts of climate change on the organizations operations.
2. Engagement with external consultants or experts specializing in climate science [...]
The question is clearly looking for processes to identify and assess risks associated with climate change. Example 1. states that "climate risk assessment tools" are relevant. The paragraph states that the Integrated Management System serves to identify risks including environmental risks. In sustainability matters, climate change and environmental topics often fall under the same umbrella. Thus, yes, the paragraph is relevant for the question addressing a certified process to manage climate risks. However, also contrary arguments can be considered. We don’t exactly know whether environmental and climate topics are viewed interchangeably. An expert may know clear differentiating factors between environmental and climate matters (e.g., not all environmental problems like water pollution affect the climate). Furthermore, the environmental management system is rather a minor note in this paragraph. Additionally, it seems that, although it is a general risk management system, the "Quality, Environmental and Occupational Health & Safety (QEHS) Management System" is rather used to achieve worldwide objectives for the company. Would you deem this relevant if it was the only information obtained for a company? And what if there are fifteen more documents that are clearly relevant? How would it be labeled then? It is possible to go to lengths and depending on which expert level or context a labeler holds. In a binary relevant/irrelevant setting, both labels would be partially wrong. The reason lies in the fact that when asking whether this document is relevant to the question, the answer is "partially right".
Appendix B Creation of Question Relevance Definition
Fig. 7 shows the prompt template for the creation of the question relevance definition. We ask the model to produce a short definition on which the model should rely. Additionally, we ask the model to produce a list of examples. This structure should align with the manner an expert implicity or explicitly approaches the annotation task of labeling relevance. A definition alone would have the shortcoming that it only incorporated generic know-how. Complementing it with examples gives the expert the flexibility to extend the meaning of the terms in exemplified form. For a demonstration of the output, see App. J.
An analyst posts a <question> about a sustainability report. Your task is to explain the <question> in the context of sustainability reporting. Please first explain the meaning of the <question>, i.e., the meaning of the question itself and the concepts mentioned. And then give a list of examples, showing what information from the sustainability report the analyst is looking for by posting this <question>.
For <the questions meaning>, please start by repeating the question in the following format:
The question "<question>" is asking for information about [...]
For the <list of example information that the question is looking for>, follow the following example in terms of format:
3. Initiatives aimed at creating new job opportunities in the green economy within the company or in the broader community.
4. Policies or practices in place to ensure that the transition to sustainability is inclusive, considering gender, race, and economic status.
Here is the question:
<question>: ""{question}""
Format your reply in the following template and keep your answer concise:
Meaning of the question: <the questions meaning>
Examples of information that the question is looking for: <list of example information that the question is looking for>"""
Figure 7: RAG Prompt Template enforcing structured output.
Appendix C Metrics Computation Details
In this project, we use Scikit-Learn (version 1.2.2) to compute AUROC, average precision scores, Brier scores, and F1 scores. We employ rank_eval (version 0.1.3) to compute nDCG and MAP scores, and Scipy.stats to compute Kendall’s $\tau$. For nDCG, relevant scores 0.5 and 1 are assigned to partially relevant and relevant documents correspondingly.
Appendix D PDF Parsing and Document Length
We use IBM deepsearch parser (Team, 2022) to parse corporate reports into chunks. For chunks shorther than 120 tokens, we concatenate them with adjacent chunks to form chunks longer than 120. Figure Fig. 8 shows the formatted chunks length distribution.
Figure 8: Distribution of chunk length after being extracted from sustainability reports and concatenation.
Appendix E Expert Annotation Process
To obtain an expert-annotated test set, we annotate the test data that contains equally 330 samples of each top-5 relevant and non-to-5 relevant documents to the 11 questions/queries, thus obtaining 660 (query, document) pairs including the corresponding expert definitions. As described in App. D, the data is obtained from real sustainability reports and split into chunks of around 150 words with the IBM deepsearch parser (Team, 2022). Table 6 shows an overview of statistical properties of the number of words in test set data.
Number of words per document
Dataset Size Mean Std Min 25% 50% 75% Max
660 150 28.5 107 131 143 162 318
Table 6: Statistical properties of the number of words in ChatReport test set data.
Then, we form a group of three expert annotators. The expert annotators comprise one graduate and one PhD student working in NLP for climate change. These two experts label the entire dataset with three labels: the document is relevant, partially relevant, or not relevant for the query including the definition. Following a simple annotation guideline:
  • Please first carefully read the provided relevance definition to understand what the question is looking for. The definition consists of a question explanation and examples of relevant information.
  • If a paragraph clearly fall into the definition of relevance, i.e., explicitly mentioned by the question explanation or examples, please annotate relevant.
  • If the paragraph is not explicitly covered by the definition but you think it somehow helps answering the question. Please annotate partially relevant.
  • Otherwise please annotate irrelevant.
Additionally, one PhD student focusing on climate change and sustainability research serves as a subject-matter meta annotator to resolve conflicts or investigate cases where both labelers arise at the label partially.
Comparing the two base annotators in the setup, we can calculate inter-annotator agreement. The Cohen’s kappa between the two labelers is 0.683 (substantial agreement). We also calculate annotators’ agreement on partial relevance. The Cohen’s Kappa turns out to be 0.129, suggesting that there are uncertainty and subjectivity associated with partial labels.
Besides the relevance, we also obtain an uncertainty label whenever there is strong disagreement (co-existence of relevance and irrelevance labels) or agreement on partial relevance (two or more annotators agree on partial relevance), the data point is labeled as uncertain. There are 103 (557) uncertain (certain) (query, document) pairs in the dataset.
Finally, the third expert annotator resolves the existing conflicts in the dataset. This results in a label distribution of Table 7. It becomes apparent that the majority of documents are not relevant while still a significant number is labeled as partially relevant and relevant.
Label Occurance
Relevant 121
Partially 65
Not Relevant 474
Table 7: Label distribution in the ChatReport test dataset.
Appendix F LLM Fine-Tuning Settings
We use the default QLoRA hyperparameter settings  [^8] https://github.com/jondurbin/qlora, namely, an effective batch size of 32, a lora r of 64, a lora alpha of 16, a warmup ratio of 0.03, a constant learning rate scheduler, a learning rate of 0.0002, an Adam beta2 of 0.999, a max gradient norm of 0.3, a LoRA dropout of 0.1, 0 weight decay, a source max length of 2048, and a target max length of 512. We use LoRA module on all linear layers. All fine-tunings last 2 epochs.
All experiments are conducted on two clusters, one with 4 V100 GPUs and the other with 4 A100 (80G) GPUs. 1 GPU hour is used per fine-tuning.
Appendix G DIRAS Prompt Template $\mathcal{P}$
Fig. 9 shows the full prompt DIRAS prompt template for the Chain-of-Thought setup. The non-CoT setup just excludes the “[Reason]: …"" part of the prompt.
You are a helpful assistant who assists human analysts in identifying useful information within sustainability reports for their analysis.
You will be provided with a <question> the analyst seeks to answer, a <paragraph> extracted from a lengthy report, and <background_information> that explains the <question>. <background_information> first explains the <question> and then raises examples to help you to better understand the <question>. Your job is to assess whether the <paragraph> is useful in answering the <question>.
<background_information>: "{background_information}"
<question>: "{question}"
<paragraph>: "{paragraph_chunk}"
Is <paragraph> helpful for answering <question>? Note that the <paragraph> can be helpful even it only addresses part of the <question> without fully answering it. Provide your best guess for this question and your confidence that the guess is correct. Reply in the following format:
[Reason]: <Reason why and how the paragraph is helpful or not helpful for answering the question. Clearly indicate your stance.>
[Guess]: <Your most likely guess, should be one of "Yes" or "No".>
[Confidence]: <Give your honest confidence score between 0.0 and 1.0 about the correctness of your guess. 0 means your previous guess is very likely to be wrong, and 1 means you are very confident about the guess.>
Figure 9: Full DIRAS Chain-of-Thought prompt for LLMs predicting relevance labels and calibrating.
Appendix H Alternative Prompts
Fig. 9, Fig. 10, and Fig. 11 show the alternative prompts with which we experimented.
{Same task description and inputs}
Is <paragraph> helpful for answering <question>? Note that the <paragraph> can be helpful even it only addresses part of the <question> without fully answering it. Provide your best guess for this question and the probability that the <paragraph> is helpful. Reply in the following format:
[Reason]: <Reason why and how the paragraph is helpful or not helpful for answering the question. Clearly indicate your stance.>
[Guess]: <Your most likely guess, should be one of "Yes" or "No".>
[Probability Helpful]: <The probability between 0.0 and 1.0 that the <paragraph> is helpful to the <question>. 0.0 is completely unhelpful and 1.0 is completely helpful.>
Figure 10: Output requirements for the alternative prompt setting $\mathcal{P}_{prob}$. Task description and input are the same as Fig. 9.
You will be provided with a <question> the analyst seeks to answer, and a <paragraph> extracted from a lengthy report. Your job is to assess whether the <paragraph> is useful in answering the <question>.
<question>: "{question}"
<paragraph>: "{paragraph_chunk}"
{Same output requirements}
Figure 11: Task description and input part for the alternative prompt setting $\mathcal{P}_{w\text{/}o\_e}$. Output requirements are the same as Fig. 9.
<|system|>
You are RankLLM, an intelligent assistant that can rank passages based on their relevancy to the query.
<|user|>
I will provide you with {num} passages, each indicated by a numerical identifier [].
Rank the passages based on their relevance to the search query: {query}.
{passages}
Search Query: {query}.
Rank the {num} passages above based on their relevance to the search query. All the passages should be included and listed using identifiers, in descending order of relevance. The output format should be [] > [], e.g., [4] > [2]. Only respond with the ranking results, do not say any word or explain.
Figure 12: We use exactly the same listwise ranking prompt as Sun et al. (2023b) and Pradeep et al. (2023). Both system and user prompts are presented in this figure.
<|system|>
You are RankLLM, an intelligent assistant that can rank passages based on their relevancy to the query.
<|user|>
I will provide you with {num} passages, each indicated by a numerical identifier [].
Rank the passages based on their relevance to the search query: {query}.
{passages}
Search Query: {query}.
Here are some background information that explains the query: {relevance_definition}
Rank the {num} passages above based on their relevance to the search query. All the passages should be included and listed using identifiers, in descending order of relevance. The output format should be [] > [], e.g., [4] > [2]. Only respond with the ranking results, do not say any word or explain.
Figure 13: Listwise prompt with an extra input of explicit definition.
Appendix I ClimRetrieve Dataset Overview
The ClimRetrieve dataset simulates the typical tasks of a sustainability analyst. The annotators examine 30 sustainability reports. We use the report-level dataset of this paper which contains 43K (query, document) pairs labelled by relevance. This dataset is very long because every report is repeated per question to have unique (query, document) pairs. We also use the much shorter, only relevant (query, document) pairs containing 595 unique samples. Since the annotators only search for relevant information, these are the gold labels. There is no active labeling of the irrelevant (query, document) pairs. Table 8 offers a comparison of the statistical properties of the word count of the documents in ClimRetrieve. They are slightly longer than the ChatReport data (compare Table 6). Table 9 shows the label distribution in the relevant-only dataset. It becomes apparent that most of the (query, document) pairs are very relevant (label=3). This aligns with the report analyst setting where information is searched for until the question can effectively be answered.
Number of words per document
Dataset Dataset Size Mean Std Min 25% 50% 75% Max
Report-Level 43.445 172 60.0 1 135 186 220 499
Only Relevant 595 199 43.8 38 177 213 228 281
Table 8: Statistical properties of the number of words in ClimRetrieve data.
Label Occurance
1 100
2 167
3 328
Table 9: Distribution of relevance labels over the relevant (query, document) pairs in ClimRetrieve.
Appendix J Creation of the Expert-Informed Relevance Definition
Fig. 14 shows the prompt for the creation process of the expert-informed relevance definitions. Following the procedure in Schimanski et al. (2024b), we make use of the text parts labeled as relevant. There exists a relevance score from 1-3 where 1 signals the least and 3 is most relevant. Similar to the base setup for the experiments in Schimanski et al. (2024b), we use the text samples with a score of 2 or higher to create the expert-informed relevance definition. We include every relevant text part as an example. Thus, the prompt includes a large set of examples per question that can be synthesized in the relevance definition. While this procedure creates an expert-informed definition, it also represents a data leakage. In a sense, the definition will retrofit with the documents it searches for. While this is a limitation, we argue that this is also the step an expert human would take when annotating. She will have fixed concepts in her head, maybe even inspired by prior search processes. Therefore, we argue that this data leakage experiment is still adequate to investigate whether an expert-informed definition helps align the search process.
Plugging the examples into the prompt results in a set of expert-informed relevance relevance definitions. When comparing these relevance definitions to the generic ones, it becomes apparent that GPT-4 already incorporated the majority of the concepts that the experts were looking for. Therefore, the adjustment of the relevance definition is visible but rather subtle. One example is displayed in App. J. While the meaning of the question remains rather static, there are nuanced differences in the examples that guide the relevance labeling.
An analyst posts a <question> about a sustainability report. Your task is to explain the <question> in the context of sustainability reporting. Please first explain the meaning of the <question>, i.e., meaning of the question itself and the concepts mentioned. And then give a list of examples, showing what information from the sustainability report the analyst is looking for by posting this <question>.
For <the questions meaning>, please start by repeating the question in the following format:
The question "<question>" is asking for information about [...]
For the <list of example information that the question is looking for>, following the following example in terms of format:
3. Initiatives aimed at creating new job opportunities in the green economy within the company or in the broader community.
4. Policies or practices in place to ensure that the transition to sustainability is inclusive, considering gender, race, and economic status.
Here is the question:
<question>: ""{question}""
Additionally, here is a <list of question-relevant example information> that an expert human labler annotated. Please keep these examples in mind when answering:
--- [BEGIN <list of question-relevant example information>]
{examples}
--- [END <list of question-relevant example information>]
Format your reply in the following template and keep your answer concise:
Meaning of the question: <the questions meaning>
Examples of information that the question is looking for: <list of example information that the question is looking for>"""
Figure 14: RAG Prompt Template enforcing structured output with the inclusion of examples.
Question Generic Definition Expert-informed Definition
Does the company provide definitions for climate change adaptation? "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly outlined what it understands by the term ""climate change adaptation"" within its sustainability reporting. This involves understanding how the company perceives the process of adjusting to actual or expected climate change and its effects, as well as how it plans to manage or mitigate harm or exploit beneficial opportunities. Examples of information that the question is looking for: 1. A clear definition of climate change adaptation as understood and applied by the company in its operations and planning. 2. Descriptions of the specific impacts of climate change that the company is adapting to, such as increased frequency of extreme weather events, rising sea levels, or changing agricultural conditions. 3. Examples of measures the company is taking to adapt to these impacts, such as investing in flood defenses, altering supply chain logistics, or developing drought-resistant crops. 4. Information on how the companys definition of climate change adaptation aligns with or differs from standard definitions provided by authoritative bodies like the Intergovernmental Panel on Climate Change (IPCC) or the United Nations Framework Convention on Climate Change (UNFCCC). 5. Details on any collaborative efforts with governments, NGOs, or other corporations to address climate change adaptation collectively. 6. Insights into how the company measures or evaluates the effectiveness of its adaptation strategies." "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly defined what it means by climate change adaptation within its sustainability reporting. This involves understanding how the company interprets climate change adaptation, including any specific strategies, initiatives, or policies it has in place to adjust to current or expected future climate change impacts. Examples of information that the question is looking for: 1. Descriptions of how the company defines climate change adaptation in the context of its operations and strategic planning. 2. Examples of specific adaptation measures the company has implemented or plans to implement, such as enhancing infrastructure resilience, diversifying water sources, or adjusting agricultural practices. 3. Information on how the companys definition of climate change adaptation aligns with or diverges from standard definitions provided by environmental organizations or regulatory bodies. 4. Details on how the company assesses and integrates climate change risks and opportunities into its investment decision-making processes, focusing on adaptation. 5. Statements on the companys involvement in partnerships or alliances aimed at promoting climate change adaptation and resilience, indicating a collaborative approach to defining and addressing adaptation needs."
Table 10: Example of a generic and expert-informed relevance definition for a question.
Appendix K MAP and nDCG Scores for Different Relevance Definitions
In the expert-informed definition experiment, we compare two settings. First, we compare the predictions on the 595 relevant-only (query, document) pairs. This is a replication of the setting in § 4.1. Since we don’t have non-relevant samples, we can only compare the nDCG. Table 11 shows the results. It becomes apparent that only for the general nDCG score, the expert-informed query rates better. For the nDCG@5, and nDCG@10, the best-performing model remains with the generic prompt. The picture turns again when widening to nDCG@15. This could be a result of the definition creation. We use examples of relevance labels 2 and 3 to create the expert-informed definition. Thus, we implicitly equalize relevance 2 and 3 in importance. This means we are likely less effective in differentiating between 2 and 3. This could explain the lower results for at lower k’s where differentiating between 2 and 3 is important v.s. the overall nDCG where differentiating between 1 and 2/3 plays a more important role.
This intuition is reinforced by the second setting, comparing the predictions on all 43K (query, document) pairs. In this setting, we also calculate the relevance for a large amount of non-relevant pairs. As Table 12 shows, the expert-informed definition now seems effective, especially when comparing MAP. MAP is agnostic to the actual degree of relevance and rather just differentiates between relevant and not relevant. Thus, the clearly higher MAP scores show that the expert-informed definition helps in differentiating between the non-relevant pairs where the definition is not meant for vs. those the definition was created with and for. This indicates that our approach is indeed sensitive to adjusting the relevance definitions.
Setting nDCG nDCG@5 nDCG@10 nDCG@15
Llama3-Askgeneric 77.23 67.60 66.18 67.57
Llama3-Tokgeneric 76.55 67.20 66.23 65.83
Llama3-Askinformed 76.52 63.24 65.69 66.39
Llama3-Tokinformed 77.41 65.95 65.06 66.91
Table 11: Comparison of using the generic and the expert-informed relevance definitions for ranking relevant only ClimRetrieve (query, document) pairs.
Setting nDCG nDCG@5 nDCG@10 nDCG@15 MAP MAP@5 MAP@10 MAP@15
Llama3-Askgeneric 29.95 18.67 21.71 23.38 26.51 17.86 21.21 22.75
Llama3-Tokgeneric 31.17 20.35 23.21 25.17 28.73 19.58 23.15 25.05
Llama3-Askinformed 30.89 19.01 22.82 24.89 29.31 20.02 23.60 25.56
Llama3-Tokinformed 32.53 21.47 24.99 26.92 32.65 22.97 27.20 28.77
Table 12: Comparison of using the generic and the expert-informed relevance definitions for ranking all ClimRetrieve (query, document) pairs.
Appendix L Embedding Fine-Tuning
We follow the official fine-tuning example [^9] https://github.com/FlagOpen/FlagEmbedding/tree/master/examples/finetune of (Chen et al., 2024) to fine-tune the embedding models. The models are fine-tuned on all annotated (query, document) pairs in ChatReport test set for 10 epochs, with a batch size of 4. Other hyperparameters are the same as the official example.
Appendix M Hand-Checking of Model Results on ClimRetrieve
The ClimRetrieve dataset presents the opportunity to investigate the output of our best-performing fine-tuned model Llama3 without CoT from different angles. In this experiment, we hand-annotate the output of our model in different categories and compare it to the ClimRetrieve annotations. As outlined in § 4, one issue of our approach is that we can only approximate the mental model of the annotator to create the relevance definitions. For the analysis, we consider the predictions of the fine-tuned Llama-3 model with the expert-informed relevance definitions created in App. J as they likely present the closest approximation to the mental model of the labelers. Furthermore, the ClimRetrieve dataset only has golden labels for the (partially) relevant documents. This has implications for our results. In this setup, we view our model as a calibrated annotator acting according to the given relevance definition – i.e., our model represents a fictional golden truth. This allows us to perform a qualitative edge case analysis on the ClimRetrieve data in different categories.
The first category is the true positive classifications (model says relevant, ClimRetrieve-human says relevant). Since they are golden annotations by the ClimRetrieve annotators and align with our model, they are not checked as they likely are error-free. A more nuanced view has to be adopted for the three categories false negatives (model says relevant, ClimRetrieve-human says not relevant), true negatives (model says not relevant, human says not ClimRetrieve-relevant), and false positives (model says not relevant, human says ClimRetrieve-relevant). To obtain a qualitative understanding, we investigate the appearances (see Table 13) and confidences (see Table 14) of the categories. For the confidences, we use the empirically best-performing model in information retrieval Llama3-Tok (see Table 2). To perform a qualitative investigation, we sample (document, query) pairs from each category and reannotate them as being "relevant", "partially relevant" or "not relevant". These reannotations do not serve to obtain a holistic view (as already done in § 3) but rather gain insights into the special cases of the model’s predictions.
First, we investigate the false negatives (model says relevant, human says not relevant). Investigating these cases is of particular interest since we know that humans have a selection bias (Thakur et al., 2021). In the analyst scenario setup of ClimRetrieve, this circumstance is aggravated by the fact that analysts may only search for information until they deem that they can answer the question at hand. Looking at Table 13, it becomes apparent that there is a very large number of false negatives. However, Table 14 shows that the confidence in the true negatives is much higher than in the false negatives. This indicates that the false negatives contain a much larger spectrum of partial relevance. This can be an initial explanation for the inflated number of false negatives.
To perform a qualitative investigation, we sample (document, query) pairs from the false negatives and perform a reannotation with a single annotator. We have 16 unique questions/queries in the ClimRetrieve dataset and sample 6 (document, query) pairs per question/query. Furthermore, we want to investigate which role model confidence plays. Thus, the (document, query) pairs are sampled to contain two samples with a Llama3-Ask confidence above 0.9, two between 0.9 and 0.7, and two below 0.7. We choose Llama3-Ask confidence because it is easier to form thresholds and approximates the confidence well enough.
Since we know that our model and the ClimRetrieve annotation inherently disagree, we focus on our reannotation. Fig. 16 shows the resulting percentage of agreement between the model and our reannotation for increasing the confidence threshold. It becomes apparent that irrespective of the threshold, the agreement between our reannotator and the model is very high. Furthermore, increasing the threshold correlates with a high model-reannotation agreement. This indicates that the model’s assignment of the "relevant" label indeed works. At the same time, it bears the question of whether the initial ClimRetrieve-human annotation is invalid. Why does it underestimate the true relevant documents by such a large magnitude? There are three avenues to explain these results. First, as already mentioned, humans may miss out on sources, suffer from selection bias, not taking duplicate information into account, and may stop annotating once they reach a sufficient level of information for a question. While these things are inherent in the analyst setting of ClimRetrieve and out of our control, we can observe that ClimRetrieve clearly misses out on some relevant information (see for instance Fig. 15). However and second, one further major reason for the inflated number of false negatives may be the change of definition we make by creating it with GPT-4. The original questions all concern the specific topic of climate change adaptation. This is a very narrow, specialized case of the general climate change domain. In our definitions, this generally looks different. Consider for instance the following example:
Meaning of the question: The question "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?" is asking for information about whether the companys stated goals or objectives for environmental sustainability or climate change mitigation are aligned with, or make reference to, established external goals or targets. These external references could include international agreements, national policies, or standards set by recognized organizations focused on climate change and sustainability.
Examples of information that the question is looking for:
1. In line with our commitment to the Net-Zero Banking Alliance (NZBA) [...]
While the question itself only addresses "climate change adaptation", the relevance definition allows for contents that are in line with "climate change mitigation". Mitigation is much broader and much more discussed in sustainability reports. The examples in the definition additionally broaden the scope. This also hints at the third avenue to explain the many false negatives: they contain a lot of partially relevant (document, query) pairs. This can also explain the significant drop in average confidence between true and false negatives in Table 14. However, the important aspect of the DIRAS pipeline is that it is consistent with the provided definitions which Fig. 16 indicates.
This investigation unveils clear edge cases for partial relevance and the mismatch of different definitions, i.e. mental models of annotators. Nonetheless, Fig. 16 reveals that, even amongst these hard examples, the approach remains consistent with the provided definitions. Although looking at edge cases, our reannotations are largely aligned with our model when overturning the non-golden irrelevant documents of ClimRetrieve.
To further get an intuition on edge cases, we turn our attention to the false positives. These represent the most complicated cases where the ClimRetrieve human annotator deemed the (document, query) pair relevant and the model arose with the judgment of not relevant. Since only 148 (document, query) pairs are in the false positives, we randomly sample 96 (document, query) pairs and reannotate them. The first two quantitative observations remain similar to the false negatives. First, Table 14 shows that the average confidence for the false positives is significantly lower than those of the true positives indicating they present edge cases. Indeed, they have the lowest average confidence among all categories with 0.8871. However, Fig. 17 again shows that with a rising confidence threshold, the agreement between the prediction and our reannotation rises. This time, the level of partial relevance is higher than for the true negatives. These two cues towards higher complexity continue in the qualitative assessment of our reannotations. We repeat to ask the question why is there a discrepancy between our model and the ClimRetrieve annotations? This time, we identify a multitude of edge cases. First, a technical edge case comes into play. ClimeRetrieve was created by matching span-labeled relevant texts with documents retrieved from the reports. A document is deemed relevant when one sentence of the span-labeled text appears in a document. However, one sentence alone cut off or in a different context may not be relevant. In 19% (18 of 96) of the annotated samples, this is the reason for a wrong label assigned. Second, the nature of partially relevant labels seems to be ambiguous. While the ClimRetrieve human annotator deemed many samples partially relevant, our reannotation overturned some of these cases in line with our model. Nonetheless, there remains a significant chunk of partially relevant samples (see Fig. 17). Third, using our model, we overturn some of the decisions of the ClimRetrieve annotators entirely. Even if they deemed it relevant, our reannotation and model suggest irrelevance given the definition. This can again be linked to the mismatch of the exact mental model of the ClimRetrieve annotators and our approximated reannotation guidelines. It has to be noted that these decisions may again be overruled by a second or third reannotator. The samples generally all have a connection with the topics in the question. Rather than presenting a gold standard, this comparison allows us to understand the fine-grained nuances of relevance, therefore reinforcing the very need of this project. Clear misannotations were only marked in 3 of 96 cases where our annotator could not identify any remote relevance.
Finally, we investigate the large chunk of the (document, query) pairs, the true negatives (model says not relevant, human says not relevant). Having a large basis of 37.409 (document, query) pairs, we again sample 6 pairs per question using two samples with a Llama3-Ask confidence above 0.9, two between 0.9 and 0.7, and two below 0.7. The true negatives are not at all comparable with the investigations before because we now sample from pairs with apriori agreement. Thus, we don’t investigate edge cases in this category.
This is also confirmed by the results (see Fig. 18). Only two samples with a Llama3-Tok confidence threshold below 0.5 were labeled as partially relevant, none were labeled as relevant. This reinforces the general results in § 3 and shows that the previous two reannotations concerned edge cases.
Collectively, these observations prove the motivation for this project in developing a nuanced stand beyond binary and even partial relevance. They show that the models function consistently with themselves - even in edge-case scenarios. This can mean something different than consistent with human annotators who cannot share their mental model.
"Golden" model with
"golden" definition
True False
Human with own mental model True
447
148
False
5.441
37.409
Table 13: Comparison of the Llama-3’s relevance prediction vs. the ClimRetrieve human annotators on ClimRetrieve’s 43K (document, query) pairs.
Labels assigned by
fine-tuned LLama3
Average Confidence
Llama3-Tok
all 0.9681
positives 0.9049
negatives 0.978
true positives 0.9575
false positives 0.8871
true negatives 0.9784
false negatives 0.9006
Table 14: Average confidence scores of Llama3-Tok for the different classification categories.
relevance definition:
Meaning of the question: The question "Has the company identified any synergies between its climate change adaptation goals and other business goals?" is asking for information about how the companys efforts to adapt to climate change are aligned with, or can complement, its other business objectives. This includes understanding if climate change initiatives also support broader strategic goals such as cost reduction, risk management, innovation, market expansion, or reputation enhancement.
Examples of information that the question is looking for:
1. Descriptions of [...]
MISSED OUT DOCUMENT:
[...] Along with these efforts, in each business segment, Sony develops and enhances risk management and business continuity plans (BCPs) from the perspective of improving risk management across supply chains, through the identification, analysis, and assessment of business continuity risks. Flood damage has grown in recent years due to the impact of climate change, prompting Sony to reassess the flood risk at its manufacturing sites in Japan and implement preventative measures that will mitigate flood damage and facilitate rapid recovery. Sony is collaborating with relevant companies and organizations, and conducts hands-on drills to address foreseeable risks, in an effort to enhance business continuity and accelerate flood recovery. Sony will continue to increase its resilience to climate change, based on its analyses and initiatives.* A global initiative in which participating corporations aim to operate on 100% renewable electricity. It is headed by an international non-governmental organization, the Climate Group, in partnership with the CDP.
Figure 15: False Negative example of clearly relevant but missed out information in ClimRetrieve.
Figure 16: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for false negatives.
Figure 17: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for false positives.
Figure 18: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for true negatives.

### v2

@@
-DIRAS: Efficient LLM-Assisted Annotation of Document Relevance in Retrieval Augmented Generation
+DIRAS: Efficient LLM Annotation of Document Relevance for Retrieval Augmented Generation
 Abstract
 Retrieval Augmented Generation (RAG) is widely employed to ground responses to queries on domain-specific documents.
-But do RAG implementations leave out important information or excessively include irrelevant information?
-To allay these concerns, it is necessary to annotate domain-specific benchmarks to evaluate information retrieval (IR) performance, as relevance definitions vary across queries and domains.
-Furthermore, such benchmarks should be cost-efficiently annotated to avoid annotation selection bias.
-In this paper, we propose DIRAS (Domain-specific Information Retrieval Annotation with Scalability), a manual-annotation-free schema that fine-tunes open-sourced LLMs to annotate relevance labels with calibrated relevance probabilities.
-Extensive evaluation shows that DIRAS fine-tuned models achieve GPT-4-level performance on annotating and ranking unseen (query, document) pairs, and is helpful for real-world RAG development.
-[^1] We will open-source all our codes, LLM generations, and human annotations.
-DIRAS: Efficient LLM-Assisted Annotation of Document Relevance
-in Retrieval Augmented Generation
-Jingwei Ni1,2thanks: Equal Contributions., Tobias Schimanski2footnotemark: , Meihong Lin4, Mrinmaya Sachan1, Elliott Ash1, Markus Leippold2,3 1ETH Zürich       2University of Zurich       3 Swiss Finance Institute (SFI) 4 University of Electronic Science and Technology of China {jingni, msachan, ashe}@ethz.ch, meihong_lin@uestc.edu.ch {tobias.schimanski, markus.leippold}@df.uzh.ch
+But do RAG implementations leave out important information when answering queries that need an integrated analysis of information (e.g., Tell me good news in the stock market today.)?
+To address these concerns, RAG developers need to annotate information retrieval (IR) data for their domain of interest, which is challenging because (1) domain-specific queries usually need nuanced definitions of relevance beyond shallow semantic relevance; and (2) human or GPT-4 annotation is costly and cannot cover all (query, document) pairs (i.e., annotation selection bias), thus harming the effectiveness in evaluating IR recall.
+To address these challenges, we propose DIRAS (Domain-specific Information Retrieval Annotation with Scalability), a manual-annotation-free schema that fine-tunes open-sourced LLMs to consider nuanced relevance definition and annotate (partial) relevance labels with calibrated relevance scores.
+Extensive evaluation shows that DIRAS enables smaller (8B) LLMs to achieve GPT-4-level performance on annotating and ranking unseen (query, document) pairs, and is helpful for real-world RAG development.
+[^1] All code, LLM generations, and human annotations in https://github.com/EdisonNi-hku/DIRAS.
+DIRAS: Efficient LLM Annotation of Document Relevance
+for Retrieval Augmented Generation
+Jingwei Ni1,2thanks: Equal Contributions., Tobias Schimanski2footnotemark: , Meihong Lin4, Mrinmaya Sachan1, Elliott Ash1, Markus Leippold2,3 1ETH Zurich       2University of Zurich       3 Swiss Finance Institute (SFI) 4 University of Electronic Science and Technology of China {jingni, msachan, ashe}@ethz.ch, meihong_lin@uestc.edu.ch {tobias.schimanski, markus.leippold}@df.uzh.ch
 1 Introduction
-RAG has become a popular paradigm for NLP applications Gao et al.
-(2024).
+RAG has become one of the most popular paradigms for NLP applications (Gao et al., 2024).
 One core phase of RAG systems is Information Retrieval (IR), which leverages cheap retrievers to filter relevant information and thus save LLM inference costs.
-Given its cost-saving nature, IR might be a performance bottleneck for RAG (Chen et al., 2023b; Gao et al., 2024).
-Both leaving out important relevant information (low recall) as well as including excessively related but irrelevant information (low precision) may lead to severe performance down-grades (Ni et al., 2023; Cuconasu et al., 2024; Niu et al., 2024; Schimanski et al., 2024a).
-However, evaluation results on general-domain benchmarks (Thakur et al., 2021) may hardly indicate the IR performance on RAG systems, as the definition of relevance varies drastically across different domains and use cases (Bailey et al., 2008).
-See App.
-A for an example where the relevance judgment may vary widely with or without domain expertise.
-Therefore, domain-specific benchmarks need to be annotated to evaluate RAG systems in different domains.
-To address this scarcity of domain-specific IR benchmarking, we propose DIRAS, a pipeline to annotate domain-specific IR data at scale (illustrated in Fig.
-1).
-Specifically, users only need to input (1) some domain-specific queries and documents and (2) definitions for what is (ir)relevant for each query.
-Then, DIRAS distills relevance prediction data from GPT-4 to fine-tune open-sourced LLMs, which can then be used to annotate relevance on a large scale without extra API cost, avoiding annotation selection bias (Thakur et al., 2021) with limited expenditure.
-Prior work distills open-sourced LLM-based rerankers using pairwise (Qin et al., 2024) or listwise (Sun et al., 2023b; Pradeep et al., 2023) as the ranking paradigm.
-However, we choose a pointwise approach (i.e., predicting a relevance score per document and then ranking according to the scores) for DIRAS, to fulfill our three Desiderata for a document relevance annotator:
-D1.
-Efficient and Effective: When benchmarking IR of RAG systems, it is important to annotate all (query, paragraph) pairs to avoid annotation selection bias, since information thoroughness is critical for many queries (e.g., overall assessment, Ni et al., 2023).
-Therefore, the pointwise method is more favorable as it largely outperforms list- or pairwise method in efficiency (Sun et al., 2023a).
-Prior work worries that pointwise ranking is efficient but not effective due to the difficulty in calibration (Sun et al., 2023a; Qin et al., 2024).
-However, this method lacks empirical investigation in prior work.
-In our work, we observe both GPT-4 and DIRAS fine-tuned LLMs achieve very competent pointwise ranking results thanks to good calibration.
-D2.
-Improved Leverage of the Relevance Definition: Relevance and partial relevance judgments might be subjective without domain expertise or explicit annotation guidelines (Bailey et al., 2008; Saracevic, 2008; Thomas et al., 2024; see also App.
+However, IR can be a performance bottleneck for RAG (Chen et al., 2023; Gao et al., 2024).
+Both leaving out important relevant information (low recall) as well as including excessively related but irrelevant information (low precision) may lead to severe decrease in performance (Ni et al., 2023; Cuconasu et al., 2024; Niu et al., 2024; Schimanski et al., 2024a).
+Furthermore, evaluation results on general-domain benchmarks (Thakur et al., 2021) may hardly indicate the IR performance on RAG systems, as the definition of relevance varies drastically across different domains and use cases (Schimanski et al., 2024b; Bailey et al., 2008).
+To address these concerns, Saad-Falcon et al.
+(2023a) propose ARES to fine-tune an in-domain LM judge to evaluate context relevance.
+Although showing effectiveness in evaluating RAG systems for QA datasets like HotpotQA (Yang et al., 2018), ARES does not address two significant Challenges that real-world RAGs are faced with:
+Figure 1: Overview of the functionality of DIRAS taking (query, relevance definition, document) triplets as input and output a binary verdict and a well-calibrated relevance score, which is sensitive to the grey-scale of partial relevance.
+C1.
+IR Recall: RAG systems are more generally purposed than QA models, where many user queries require an integrative information analysis.
+For example, Write an overview of major events in World War 2; or What is good news for the stock market today?
+For such integrative queries, a good IR recall is necessary for comprehensive responses.
+Saad-Falcon et al.’s (2023a) approach evaluates the relevance of retrieved contexts (precision), but is agnostic to other important information that the RAG retriever might leave out (recall).
+Besides, current RAG literature (Yan et al., 2024; Wang et al., 2024) mostly relies on QA datasets (Joshi et al., 2017; Yang et al., 2018; Dinan et al., 2019; Trivedi et al., 2022) for evaluation, where the questions are less integrative and can mostly be answered by specific facts from one or few sources.
+For such questions, IR precision is more important than recall.
+As a result, the challenges of IR recall for integrative queries are heavily under-explored.
+C2.
+Relevance Definitions and Partial Relevance: To thoroughly gather relevant information for integrative queries, the IR model should go beyond shallow semantic relationships and consider domain-specific relevance definitions.
+Furthermore, domain-specific requirements and subjectivity in IR annotation create a rich gray scale of partial relevance between relevant and irrelevant (Bailey et al., 2008; Saracevic, 2008; Thomas et al., 2024; also see App.
 A).
-Therefore, DIRAS explicitly puts relevance definition into relevance prediction prompts to achieve more objective and consistent results.
-Compared to the list- or pairwise method, the pointwise method is provided with only one document each time.
-Thus, it may analyze the document along the relevance definition in better detail, especially with CoT prompting (Wei et al., 2022).
-D3.
-Richer Predictions for RAG Requirements: Listwise or pairwise ranking algorithms only predict a relative rank for documents.
-With only the rank, retrieving the same number of documents (top-k) for all questions is suboptimal as it is likely that different questions have different amounts of relevant information (details in § 5).
-In contrast, the pointwise method predicts not only ranks but also binary relevance and calibrated relevance probabilities.
-Both binary labels and relevance scores allow RAG systems to retrieve the actual amount of relevant information to a question.
-Furthermore, the calibrated relevance probability is also helpful in automatic annotation (Ni et al., 2024), indicating which annotations are partially relevant and/or check-worthy.
-To evaluate DIRAS, we conduct experiments with two datasets.
-First, we annotate a high-quality dataset based on ChatReport [^2] https://reports.chatclimate.ai/ (Ni et al., 2023), a real-world RAG application analyzing lengthy corporate reports.
-This dataset also incorporates the ideas of partial relevance and labeling uncertainty.
-ChatReport is representative of RAG applications that are sensitive to retrieval results, as thoroughly analyzing disclosed information is crucial for assessing what is under-addressed in the reports.
-Evaluation on ChatReport data shows that the pointwise ranking of DIRAS achieves very satisfactory performance, even superior to the widely-adopted listwise method (Pradeep et al., 2023) (§ 3.1).
-The best DIRAS fine-tuned model also achieves GPT-4-level performance in terms of relevance ranking and calibration (§ 3.2).
-Second, we use a DIRAS fine-tuned model to re-annotate all (query, document) combinations (43k in total) in ClimRetrieve (Schimanski et al., 2024b), a real-world record of experts’ information seeking workflow.
-Experiments show that DIRAS fine-tuned models successfully understand fine-grained degree of relevance (§ 4.1) and enhance their performance through improved relevance definitions (§ 4.2).
-Furthermore, it mitigates IR annotation bias by identifying information ignored by experts (§ 4.3), and benchmarks IR algorithms’ target domain performance upon all 43k (query, document) pairs (§ 4.4).
-Finally, we propose recommendations for future RAG designs based on our takeaways (§ 5).
+However, partial relevance is neglected entirely in RAG context relevance evaluation (Saad-Falcon et al., 2023a; Es et al., 2024).
+As a combined solution for these challenges, we propose DIRAS, a framework for efficient and effective relevance annotation.
+To address C1, DIRAS distills relevance annotation ability from SOTA generic teacher LLMs to small student LLMs, which can cost-efficiently annotate broad (query, document) pairs for IR recall evaluation.
+For better efficiency, student LLMs conduct pointwise annotation – annotating (query, document) pairs one-by-one – which is under-explored in related work (Sun et al., 2023a; Qin et al., 2024) but achieves good performance for relevance annotation.
+To address C2, DIRAS student LLMs are trained to comprehend nuanced relevance definitions, thus handling queries with various requirements.
+The student LLMs annotate binary relevance labels with well-calibrated relevance scores.
+Thus, the relevance scores can be used for relevance ranking and to calibrate the annotation accuracy (Ni et al., 2024).
+Thereby, these continuous relevance scores also address the grayscale of partial relevance (see Fig.
+1 for an illustration of DIRAS functionality).
+We evaluate DIRAS in three steps.
+First, we annotate ChatReportRetrieve to evaluate the design decisions for making DIRAS models optimized relevance annotators (§ 3.1).
+The evaluation shows that the fine-tuned student ($\leq$ 8B) LLMs effectively understand nuanced relevance definitions – achieving GPT-4-level performance (§ 3.3).
+Second, we showcase how DIRAS assists in real-life IR annotation by re-annotating ClimRetrieve (Schimanski et al., 2024b).
+Results show that DIRAS student LLMs can effectively capture partial relevance, leverage improved relevance definitions, mitigate annotation selection bias, and annotate benchmarking datasets for IR algorithms (§ 4.1).
+Third, we re-annotate document relevance for general QA and RAG datasets, showing DIRAS’s potential in generic domains (§ 4.2).
 Collectively, our contributions include:
   1. 1.
-    We propose DIRAS, a framework tuning open-sourced LLMs into efficient and effective IR annotators, taking domain expertise into account.
+    We propose DIRAS, a framework fine-tuning open-sourced LLMs into efficient and effective IR annotators, taking domain expertise into account.
   2. 2.
-    We annotate a high-quality IR benchmark based on ChatReport with (partial) relevance labels and uncertainty labels, based on explicit relevance definitions.
+    We annotate ChatReportRetrieve, the first IR benchmark addressing integrative queries in RAG, and providing explicit relevance definition as annotation guidelines.
   3. 3.
-    We compare DIRAS fine-tuned models with a real-world information-seeking workflow by experts, showing the model accurately understands granular relevance definitions and helps mitigate annotation selection bias.
-Figure 1: DIRAS pipeline. RAG-specific queries, documents as inputs; calibrated small-LLM annotators as outputs.
-2 DIRAS
-2.1 DIRAS Pipeline
-To address the outlined desiderata (D1, D2, D3) in § 1, we design the DIRAS pipeline (illustrated in Fig.
-1).
-This pipeline comprises the training data creation and fine-tuning of small LLMs to the calibrated annotators.
-Training Data Creation: We create the training data from domain-specific sources like reports.
-As a result, the data will be composed of a set of domain-specific (query, document) pairs.
-Each query comprises a question and a definition indicating what is relevant or irrelevant to the question.
-The relevance definition can be designed by human experts or generated by LLMs following App.
+    We showcase how to apply DIRAS in real-world IR annotation by re-annotating ClimRetrieve and general QA datasets.
+Figure 2: DIRAS pipeline. Domain-specific queries, and documents as inputs; calibrated student LLM annotators as outputs.
+2 DIRAS Pipeline
+The DIRAS pipeline is illustrated in Fig.
+2.
+It comprises three steps: sampling a subset of (query, document) pairs for training data creation, obtaining relevance definitions for queries, and distilling relevance annotation ability from teacher to student LLMs.
+Sampling (Query, Document) Pairs: DIRAS takes domain-specific queries and documents as input.
+To sample representative (query, document) pairs as training data, it first ranks documents for each query using a small dense retriever.
+Then, it samples an equal number of documents within and outside of top-k (a pre-defined hyperparameter) to obtain representative documents for each query.
+While sampling in top-k aims at covering some relevant documents, sampling outside of top-k ensures covering the broader distribution of (query, document) pairs.
+Obtaining Relevance Definitions: Each query in the sampled (query, document) pairs needs to be accompanied by an explicit definition indicating what is relevant or irrelevant to the question.
+The relevance definition can be generated by human experts, LLMs, or in a collaboration of both.
+In our experiments, we find GPT-4 generates suitable relevance definitions using the prompt in App.
 B.
-To obtain the documents for each question/query, we employ a sampling strategy using top-k relevant documents ranked by a small dense retriever.
-We set a top-k threshold and sample an equal number of documents within and outside of the threshold for each question.
-While sampling in top-k aims at covering some relevant documents, sampling outside of top-k ensures covering the broader distribution of (query, document) pairs.
-Open-Sourced LLM Fine-Tuning: Once we sampled the domain-specific (query, document) pairs, we create instruction fine-tuning (IFT) data with a SOTA generic LLM $\mathcal{M}$ and the prompt template $\mathcal{P}$ (see e.g., Fig.
-2).
-The prompt template incorporates a (query, document) pair and an instruction to predict a binary relevance label and its confidence score.
-Finally, the distilled IFT data is used to fine-tune open-sourced LLMs $\mathcal{M}_{o}$ (“o” stands for “open”) to conduct binary relevance prediction with confidence.
-2.2 Evaluation Metrics
-Both $\mathcal{M}$’s and $\mathcal{M}_{o}$’s results should be evaluated against two types of human labels.
-(1) Relevance labels: whether a document is helpful for answering a query or not, disagreement between different annotators needs to be resolved to obtain the final relevance label.
-(2) Uncertainty labels: an annotation is uncertain, if the annotators have a strong disagreement, or the majority of them agree that the (query, document) is partially relevant.
-We use the following metrics for evaluation:
-Binary Relevance: We compute the F1 Score of models’ binary relevance prediction using relevance labels.
-For RAG systems, binary relevance labels are important for deciding which documents should be passed to LLMs.
-Calibration: Both $\mathcal{M}$ and $\mathcal{M}_{o}$ give confidence scores, which should calibrate the binary accuracy to indicate whether the prediction is trustworthy.
-We use Expected Calibration Error (ECE), Brier Score, and AUROC to measure calibration performance, following Kadavath et al.
-(2022) and Tian et al.
-(2023).
-Information Retrieval: The confidence scores also give a calibrated relevance probability which can be used to rank documents for each query.
-To directly evaluate the ranking performance, we measure nDCG and MAP upon relevance labels.
-Uncertainty: If the models understand the difficulty and uncertainty caused by partial relevance, they should have lower confidence scores on samples that humans found to be uncertain.
-Thus we compute average precision (AP) scores between confidence and uncertainty labels.
-We chose AP because it is threshold-free by summarizing F1 scores of all thresholds of relevance scores.
-Details of computing all metrics are in App.
-C.
-3 Experiments on ChatReport
-To the best of our knowledge, there is no existing IR dataset that (1) accompanies each query with domain-specific relevance definition; and (2) provides annotations of human uncertainty caused by partial relevance.
-Therefore, we annotate a dataset based on ChatReport (Ni et al., 2023) to fulfill evaluation purposes in § 2.2.
-ChatReport is an online RAG application for corporate climate report analyses and Q&A about the reports.
-[^3] See https://reports.chatclimate.ai/.
-Data Sampling: We sample 31 questions and 80 reports from this application.
-Climate reports are sampled randomly from openly accessible user submissions [^4] See https://github.com/EdisonNi-hku/chatreport.
-(PDF parsing details in App.
-D).
-The questions are strategically sampled to ensure representativeness and diversity.
-Specifically, 11 questions are the core questions used in ChatReport, which cover essential topics of sustainability disclosure.
-20 questions are selected from users’ customized questions posed to the ChatReport tool.
-Finally, we prompt GPT-4 to draft relevance definitions for all questions (see App.
+Distilling Relevance Annotations from Teacher to Student: With the sampled (query, definition, document) triplets, DIRAS creates relevance-annotation data with a SOTA generic teacher LLM $\mathcal{M}_{t}$ and the prompt template $\mathcal{P}$ (illustrated in Fig.
+3).
+Finally, the created data is used to fine-tune student LLMs $\mathcal{M}_{s}$, which will be used to conduct broad binary relevance annotation with confidence scores for calibration (Tian et al., 2023).
+3 Optimizing Relevance Annotation
+To train a highly performant student LLM $\mathcal{M}_{s}$ for relevance annotation, we proceed in three steps.
+First, we annotate a novel task-specific dataset for our evaluation, ChatReportRetrieve (§ 3.1).
+Second, we assess design choices and related work baselines to find the best-performing strategy for our relevance-annotation data creation with the teacher model $\mathcal{M}_{t}$ (§ 3.2).
+Third, we optimize the fine-tuning of student LLMs $\mathcal{M}_{s}$ by investigating implementation variants (§ 3.3).
+3.1 ChatReportRetrieve
+To evaluate the teacher LLMs’ ($\mathcal{M}_{t}$) and student LLMs’ ($\mathcal{M}_{s}$) comprehension of nuanced relevance definitions, we need to provide them with the same annotation guidelines (i.e., relevance definitions) and compare their annotation performance.
+To the best of our knowledge, there is no existing IR dataset that provides a nuanced relevance definition for each query.
+Hence, we annotate ChatReportRetrieve for our evaluation.
+Data and annotation guideline preparation: ChatReportRetrieve is based on real-user integrative queries from ChatReport – a chat tool for answering climate-related questions based on corporate reports [^2] https://reports.chatclimate.ai/ (Ni et al., 2023).
+We sample a wide range of climate reports and representative user queries about the reports to construct ChatReportRetrieve.
+Then, we conduct a train-test split, making sure no test set report or query is seen in the training set.
+Finally, we prompt GPT-4 to draft relevance definitions for all queries.
+GPT-4 drafted relevance definitions show a good understanding of the climate disclosure domain, according to a domain expert’s feedback.
+See Appendices for details of data preprocessing (App.
+D) and relevance definition generation (App.
 B).
-Train-Test Split: We split the questions into 11 for testing and 20 for training.
-Similarly, we split the reports into 30 for testing and 50 for training.
-This ensures the evaluation on unseen questions and reports.
-For each question, we randomly sample 60 documents – 30 each from in the top-5 and outside the top-5 (using OpenAI text-embedding-3-small as the dense retriever).
-Ultimately, (query, document) pairs in training split are used to create training data with relevance label and confidence score predictions (details in § 3.1).
-Data points in the test split are passed to human annotation.
 Test Data Annotation: We leverage relevance definitions as the annotation guidelines.
-A document is relevant if and only if it fulfills the definition.
+If and only if a document addresses the relevance definition, it is deemed as (partially) relevant.
+We explicitly account for partial relevance when the document addresses the periphery of the definition.
 The data labeling process follows two steps.
@@
 E.
+Evaluation Metrics: LLM predictions contain a binary relevance annotation and a confidence score.
+They will be evaluated against relevance or uncertainty of ChatReportRetrieve labels on four dimensions: (1) Binary Relevance: We compute the F1 Score of models’ binary relevance prediction using relevance labels.
+Binary relevance labels are important for deciding which documents should be passed to LLMs.
+(2) Calibration: Confidence scores should calibrate the binary accuracy to indicate annotation quality.
+We use Expected Calibration Error (ECE), Brier Score, and AUROC to measure calibration performance, following Kadavath et al.
+(2022) and Tian et al.
+(2023).
+(3) Information Retrieval: The confidence scores also give a calibrated relevance probability which can be used to rank documents for each query.
+To directly evaluate the ranking performance, we measure nDCG and MAP upon relevance labels.
+(4) Uncertainty: If the models understand the difficulty caused by partial relevance, they should have lower confidence scores on samples that humans found uncertain.
+Thus we compute average precision (AP) scores between confidence and uncertainty labels.
+Details of computing all metrics are in App.
+C.
 Prompt: <question>: What is the firm’s Scope 3 emission?
@@
 Provide your best guess, and confidence score from 0 to 1.
-Teacher LLM $\mathcal{M}$: [Reason]: {Reason why the paragraph is (un)helpful.} [Guess]: {Yes or No.} [Confidence]: {confidence score between 0.0 and 1.0.}
-Figure 2: Our prompt template $\mathcal{P}$ for distilling training data from $\mathcal{M}$. It is shortened for presentation. Full $\mathcal{P}$ is in Fig. 9.
-3.1 How to Distill Training Data?
+Teacher LLM $\mathcal{M}_{t}$: [Reason]: {Reason why the paragraph is (un)helpful.} [Guess]: {Yes or No.} [Confidence]: {confidence score between 0.0 and 1.0.}
+Figure 3: Our prompt template $\mathcal{P}$ for distilling training data from $\mathcal{M}_{t}$.
+“[Reason]” is only used in the CoT setup.
+It is shortened for presentation.
+Full $\mathcal{P}$ is in App.
+Fig.
+9.
+3.2 Optimizing the Training Data Creation
 Setting Unc. Bin. Cal. Info. Avg.
@@
 List-20/10-D - - - 82.54 -
-Point-Ask 39.27 84.07 94.41 87.57 76.33
-Point-Ask-Prob-D 44.74 84.52 93.72 88.39 77.84
-Point-Tok-D 28.83 86.32 93.31 80.90 72.52
-Point-Ask-D 54.01 86.32 94.41 88.48 80.80
-Table 1: GPT-4’s performance on ChatReport test set with different ranking methods (Point- or Listwise), with/without relevance definition (D), and calibration method (Ask or Tok).
-Unc.
-denotes the Average Precision (AP) of predicting uncertainty; Bin.
-denotes F1 score of binary relevance prediction; Cal.
-is the average of AUROC, ECE, and Brier Score; Info.
-is the average of nDCG and MAP; Avg.
-denotes the average of all metrics.
-Best scores of each column are bolded.
-To train better open-sourced LLMs $\mathcal{M}_{o}$, it is crucial to distill high-quality training data from teacher LLM $\mathcal{M}$.
-Specifically, we compare the following three implementation choices:
+RAGAs - 68.15 - 37.13 -
+ARES-0-Shot 25.48 52.63 79.35 77.67 58.79
+ARES-2-Shot 17.38 3.85 63.49 44.97 32.42
+ARES-4-Shot 18.16 5.13 69.51 49.68 35.62
+ARES-8-Shot 16.99 28.81 66.77 48.75 40.33
+ARES-16-Shot 20.65 24.20 64.76 44.85 38.61
+Point-Ask 39.27 84.07 90.59 87.57 75.37
+Point-Ask-Prob-D 44.74 84.52 91.31 88.39 77.24
+Point-Tok-D 28.83 86.32 84.48 80.90 70.53
+Point-Ask-D 54.01 86.32 91.10 88.48 80.00
+Table 1: GPT-4’s performance on ChatReportRetrieve test set with different ranking methods (Point- or Listwise), RAGAs and ARES-few-shot, with/without relevance definition (D), and calibration method (Ask or Tok).
+Bin., Cal., Info., and Unc.
+stand for evaluation dimensions in § 3.1.
+We aim to train a highly performant student LLM $\mathcal{M}_{s}$ for relevance annotation.
+Thus, it is crucial to identify best-performing implementation choices that can be used to distill high-quality training data from the teacher LLM $\mathcal{M}_{t}$.
+Specifically, we compare the following four implementation choices:
+ARES few-shot vs.
+relevance definitions: ARES (Saad-Falcon et al., 2023a) and DIRAS have different strategies to create target-domain training data: the former uses few-shot ICL [^3] Original ARES uses few-shot ICL to create synthetic queries instead of relevance labels, which is not applicable for ChatReportRetrieve.
+Thus we use their ChatGPT prompt for relevance judgement, while replacing GPT-3.5 with GPT-4.
+while the latter uses relevance definitions.
+We also include RAGAs (Es et al., 2024) relevance judgement as a baseline.
 Pointwise vs.
-Listwise: The listwise method is popular in ranking data distillation given its moderate cost and good performance (Sun et al., 2023b; Pradeep et al., 2023).
+Listwise: The listwise method is popular in ranking data creation given its moderate cost and good performance (Sun et al., 2023b; Pradeep et al., 2023).
 However, the more efficient pointwise method is under-explored in prior work – majorly due to the concern about poor calibration (Sun et al., 2023a; Qin et al., 2024).
@@
 With vs.
-without relevance definition: As ChatReport test data is annotated based on the relevance definition, performance should increase if the model correctly takes the in-context relevance definition into consideration.
+without relevance definition: As ChatReportRetrieve test data is annotated based on the relevance definition, performance should increase if the model correctly takes the in-context relevance definition into consideration.
 Following the takeaways of Thomas et al.
-(2024), we design the prompt $\mathcal{P}$ for the pointwise method with detailed task/role description, relevance definition and CoT prompting (see Fig.
+(2024), we design the prompt $\mathcal{P}$ for the pointwise method, relevance definition and CoT prompting (see Fig.
 3 and Fig.
@@
 text-embedding-3-small is used for listwise methods’ initial ranking.
-The results are shown in Table 1.
-We observe that: (1) With the proper calibration method (Ask), the pointwise method outperforms the listwise method.
+ARES and RAGAs settings are from the original papers.
+For few-shot ICL, we keep relevant/irrelevant samples balanced to avoid bias.
+Takeaways: Results in Table 1 show that: (1) Few-shot ICL fails to teach domain-specific relevance.
+The ICL illustrations (even balanced) seem to bias GPT-4 to underperform the zero-shot setting.
+(2) With the proper calibration method (Ask), the pointwise method outperforms the listwise method.
 (3) The listwise method is sensitive to window size, while the pointwise method gives more consistent performance across prompts.
@@
 Thus we choose pointwise to be our distillation strategy.
-3.2 Open-Sourced LLM Fine-Tuning
+3.3 Optimizing DIRAS student LLMs
 Setting Unc. Bin. Cal. Info. Avg.
-Small-embed - - 83.63 66.34 -
-Large-embed - - 84.07 69.36 -
-BGE-Gemma - - 78.21 68.47 -
-GPT-3.5 29.71 45.27 87.79 74.16 59.23
-GPT-4 54.01 86.32 94.41 88.48 80.80
-Llama3-CoT-Ask 36.57 76.58 93.32 86.15 73.16
-Llama3-CoT-Tok 41.74 76.58 90.82 85.96 73.78
-Llama3-Ask 40.18 82.11 94.04 86.02 75.59
-Llama3-Tok 41.60 82.11 94.41 89.19 76.83
-Phi3-CoT-Ask 36.08 72.95 92.35 80.56 70.48
-Phi3-CoT-Tok 35.49 72.95 88.15 80.64 69.31
-Phi3-Ask 32.30 73.23 91.16 80.05 69.18
-Phi3-Tok 38.00 73.23 92.05 86.94 72.55
-Gemma-CoT-Ask 31.60 72.38 91.14 81.39 69.13
-Gemma-CoT-Tok 39.03 72.38 88.58 80.33 70.08
-Gemma-Ask 25.74 67.13 88.27 77.43 64.64
-Gemma-Tok 50.72 67.13 92.44 81.17 72.87
-Table 2: Comparison between the fine-tuned $\mathcal{M}_{o}$ and different baselines on ChatReport test data.
-Unc.
-denotes the Average Precision (AP) of predicting uncertainty; Bin.
-denotes F1 score of binary relevance prediction; Cal.
-is the average of AUROC, ECE, and Brier Score; Info.
-is the average of nDCG and MAP; Avg.
-denotes the average of all metrics.
-The best scores are bolded and the second bests are underlined.$\dagger$ denotes the best score achieved by each LLM architecture.
-DIRAS data fine-tuned models $\mathcal{M}_{o}$ will be used to predict all (query, document) combinations to mitigate annotation selection bias (Thakur et al., 2021).
-Therefore, shorter generations from $\mathcal{M}_{o}$ (e.g., without CoT) are favored as the inference cost for transformers increases quadratically with generation length.
-Additionally, the choice of calibration method matters for LLMs (Tian et al., 2023).
-To explore these aspects, we fine-tune $\mathcal{M}_{o}$ in four settings: $\mathcal{M}_{o}$-CoT-Ask, $\mathcal{M}_{o}$-CoT-Tok, $\mathcal{M}_{o}$-Ask, $\mathcal{M}_{o}$-Tok, where CoT means $\mathcal{M}_{o}$ is tuned to generate [Reason], [Guess], and [Confidence]; without CoT denotes $\mathcal{M}_{o}$ is tuned to only generate [Guess] and [Confidence]; “Ask” means the result is calibrated by the generated confidence score in [Confidence] field; and “Tok” means we take the token-level probability of “Yes/No” after “[Guess]:” as the confidence score for calibration.
+Small-embed - - - 66.34 -
+Large-embed - - - 69.36 -
+BGE-Gemma - - - 68.47 -
+GPT-3.5 29.71 45.27 85.46 74.16 58.65
+GPT-4 54.01 86.32 91.10 88.48 80.00
+Llama3-CoT-Ask 36.57 76.58 89.30 86.15 72.15
+Llama3-CoT-Tok 41.74 76.58 86.61 85.96 72.72
+Llama3-Ask 40.18 82.11 90.14 86.02 74.61
+Llama3-Tok 41.60 82.11 91.35 89.19 76.06
+Phi3-CoT-Ask 36.08 72.95 88.76 80.56 69.59
+Phi3-CoT-Tok 35.49 72.95 84.20 80.64 68.32
+Phi3-Ask 32.30 73.23 85.56 80.05 67.79
+Phi3-Tok 38.00 73.23 89.52 86.94 71.92
+Gemma-CoT-Ask 31.60 72.38 86.38 81.39 67.94
+Gemma-CoT-Tok 39.03 72.38 83.49 80.33 68.81
+Gemma-Ask 25.74 67.13 81.80 77.43 63.02
+Gemma-Tok 50.72 67.13 90.07 81.17 72.27
+Table 2: Comparison between the fine-tuned student $\mathcal{M}_{s}$ and different baselines on ChatReportRetrieve test data.
+The best scores are bolded and the second bests are underlined.$\dagger$ denotes the best score achieved by each backbone LLM.
+DIRAS student LLMs $\mathcal{M}_{s}$ will be used to annotate all (query, document) combinations.
+Two methodological choices might influence the quality of fine-tuned student LLMs.
+First, we explore the role of Chain-of-Thought (CoT) reasoning.
+Second, we investigate the choice of calibration method (Tian et al., 2023).
+To explore these aspects, we fine-tune $\mathcal{M}_{s}$ in four settings: $\mathcal{M}_{s}$-CoT-Ask, $\mathcal{M}_{s}$-CoT-Tok, $\mathcal{M}_{s}$-Ask, $\mathcal{M}_{s}$-Tok, where CoT means $\mathcal{M}_{s}$ is tuned to generate [Reason], [Guess], and [Confidence]; without CoT denotes $\mathcal{M}_{s}$ is tuned to only generate [Guess] and [Confidence]; “Ask” means the result is calibrated by the generated confidence score in [Confidence] field; and “Tok” means we take the token-level probability of “Yes/No” after “[Guess]:” as the confidence score for calibration.
 The prompt in Fig.
@@
 F).
-We compare these fine-tuned models with baselines including GPT-3.5 and GPT-4 using prompt $\mathcal{P}$; the OpenAI embedding models text-embedding-3-small, and text-embedding-3-large; and BGE Gemma reranker [^5] https://huggingface.co/BAAI/bge-reranker-v2-gemma, a popular LLM-based reranker for general domain.
+We compare these fine-tuned student models with baselines including GPT-3.5 and GPT-4 using prompt $\mathcal{P}$; the OpenAI embedding models text-embedding-3-small, and text-embedding-3-large; and BGE Gemma reranker [^4] https://huggingface.co/BAAI/bge-reranker-v2-gemma, a popular LLM-based reranker for general domain.
 As Fig.
 4 shows, fine-tuning improves original models in all settings.
-Furthermore, Table 2 shows the results of all fine-tuned models in comparison to all baselines.
-We observe that $\mathcal{M}_{o}$-Tok outperforms other settings for all LLM architectures.
+Furthermore, Table 2 shows the results of all fine-tuned student models in comparison to all baselines.
+We observe that $\mathcal{M}_{s}$-Tok outperforms other settings for all LLM architectures.
 The best setting Llama3-Tok achieves GPT-4 level performance in calibration and IR on unseen questions and reports.
@@
 CoT sometimes leads to a limited increase when asking for calibration (Ask), but constantly results in a performance drop when calibrated with token-level probability (Tok).
-Therefore, $\mathcal{M}_{o}$ should be fine-tuned without CoT for performance and inference efficiency.
+Therefore, $\mathcal{M}_{s}$ should be fine-tuned without CoT for inference efficiency.
 Moreover, Tok rarely underperforms Ask, different from Tian et al.
@@
 Figure 4: Shaded bars denote the performance of original models. Colored bars denote the improvement brought by fine-tuning.
-4 Experiments on ClimRetrieve
-In this section, we showcase how the DIRAS fine-tuned $\mathcal{M}_{o}$ can assist the IR annotation in a real-world setting, leveraging the ClimRetrieve dataset (Schimanski et al., 2024b).
-This dataset records analysts’ real-life procedure of sustainability report analyses: raise questions about sustainability reports and then go through these reports to find relevant information to answer their questions.
-The dataset comprises 43k (query, document) pairs, among which 595 pairs are annotated as relevant.
-Relevant documents are annotated with a relevance score from 1 to 3 (1 translates to partially relevant and 2/3 means relevant).
-App.
-I gives a detailed overview of the dataset.
-Compared to ChatReport data, ClimRetrieve is a more challenging and realistic setting because: (1) For ChatReport data, we draft explicit relevance definitions and annotate relevance dependent on them.
-But for ClimRetrieve, the analysts’ mental model for what is (ir)relevant for their posed questions is unknown to us.
-(2) ClimRetrieve records human analysts’ real-life workflow of reading the full reports and searching for relevant information, which differs from ChatReport data where annotators are presented with (query, document) pairs as separate data points.
-Thus, ClimRetrieve only has gold labels for relevant (query, document) pairs.
+4 Real-World Applications
+Having established and benchmarked the design choices of DIRAS student LLMs, we now investigate their real-world application capabilities.
+First, we showcase how the DIRAS $\mathcal{M}_{s}$ can assist IR annotation in a real-world setting, leveraging the ClimRetrieve dataset (Schimanski et al., 2024b) (§ 4.1).
+Second, we show the DIRAS pipeline is also applicable for general domain QA (§ 4.2).
+4.1 Applying to ClimRetrieve
+ClimRetrieve (Schimanski et al., 2024b) records human analysts’ real-life workflow of reading full reports, searching for relevant information, and annotating useful information for climate-related questions with relevance scores 1-3.
+ClimRetrieve contains 43K (query, document) pairs (8K unique documents but each document in a report is multiplied by the amount of analyzed queries per report) out of which 595 are gold labels for relevant (query, document) pairs.
 Other not annotated (query, document) combinations might be either irrelevant or a part of annotation selection bias – a widely existing problem in IR annotation (Thakur et al., 2021).
-This challenging and realistic setting of ClimRetrieve allows us to investigate the following research questions regarding DIRAS fine-tuned $\mathcal{M}_{o}$: (1) RQ1: Can $\mathcal{M}_{o}$ understand fine-grained differences in degree of relevance?
-(2) RQ2: Can we improve $\mathcal{M}_{o}$’s performance through improving relevance definitions?
-(3) RQ3: Can $\mathcal{M}_{o}$ assists in mitigating annotation selection bias (Thakur et al., 2021)?
-(4) RQ4: Can $\mathcal{M}_{o}$’s predictions help benchmarking IR algorithms?
-We use the best $\mathcal{M}_{o}$ in § 3 (Llama-3 without CoT) to study these RQs.
-4.1 RQ1: Understanding Fine-Grained Relevance Levels
+To succeed on this dataset, the IR model needs to (1) capture the analysts’ mental model about useful information (i.e., relevance definition), and (2) understand fine-grained differences in degree of relevance (score 1-3).
+Since ClimRetrieve is still in the climate domain, we re-annotate its data with the best $\mathcal{M}_{s}$ in § 3 (Llama3-Tok), and explore whether $\mathcal{M}_{s}$’s annotations can (1) RQ1: reflect fine-grained differences in degree of relevance; (2) RQ2: be improved through refining relevance definitions; (3) RQ3: mitigate annotation selection bias for better IR recall evaluation; and (4) RQ4: benchmark and select IR algorithms.
 Setting nDCG nDCG@5 nDCG@10 nDCG@15
@@
 Table 3: Performance on ranking the relevant (query, document) pairs in ClimRetrieve.
-We first evaluate fine-tuned Llama-3 on 595 gold labels of ClimRetrieve to verify whether it can effectively recover analysts’ ranking for relevant content by understanding which documents are more helpful than others.
+RQ1: Reflecting Fine-Grained Relevance Levels.
+We first evaluate Llama3-Tok’s annotation on 595 gold labels of ClimRetrieve to verify whether it can effectively recover analysts’ ranking for relevant content by understanding which documents are more helpful than others.
 Relevance definitions are drafted with GPT-4 with the same procedure as § 3.
@@
 Importantly, all ClimRetrieve annotations are to some degree relevant, so improvement over the random baseline is challenging as the system needs to understand the trivial different degrees of relevance.
-Table 3 presents different systems’ performance.
-There is a clear trend of outperformance of the fine-tuned Llama-3 models in this challenging setting.
-They also exceed the random baseline by a significant margin, indicating the model correctly understands the fine-grained levels of relevance.
-4.2 RQ2: Improving Performance through Improving Definitions
+Table 3 presents different systems’ performance. There is a clear trend of outperformance of the fine-tuned Llama-3 models in this challenging setting.
 Setting nDCG MAP
@@
 […]
-Table 4: Comparison of using the generic and the expert-informed relevance definitions for ranking all ClimRetrieve (query, document) pairs.
-We then test whether we can improve $\mathcal{M}_{o}$’s performance by improving the relevance definition presented in prompts.
-In § 3 and § 4.1, we use GPT-4 to draft the relevance definitions.
-To investigate the role of the definitions, we compare two setups: (1) The generic relevance definition: the definition drafted by GPT-4.
-(2) The informed relevance definition: we use the labeled texts in the ClimRetrieve dataset as examples to create artificial expert- textitinformed relevance definitions.
-This way, we simulate the inclusion of expert knowledge to improve the relevance definition (see App.
+Table 4: Comparison of using the generic and the improved relevance definitions for ranking all ClimRetrieve (query, document) pairs.
+RQ2: Improving Performance through Improving Definitions.
+So far, we used GPT-4 to draft the relevance definitions.
+To investigate the effect of improved definitions, we compare two setups: (1) The generic relevance definition: the definition drafted by GPT-4.
+(2) The improved relevance definition: The only way to improve the definition is to align it closer to ClimRetrieve annotators’ mental model of document relevance.
+We achieve this by adding relevant text samples to the prompt for generating the definition with GPT-4 (see App.
+I for details).
+After creating the improved definitions, we repeat predicting the relevance scores with Llama3-Tok.
+Since we involve examples with various relevance scores to improve relevance definitions, these definitions might not help distinguish the granular level of partial relevance.
+However, the improved definitions might especially help to distinguish relevant documents from irrelevant ones.
+Calculating the nDCG and MAP score for all 43K (query, document) pairs, we find evidence for this notion (see Table 4).
+Thus, the inclusion of improved definitions seems to enhance the performance (see App.
 J for details).
-After creating the definitions, we repeat predicting the relevance scores with the DIRAS fine-tuned Llama-3.
-First, we repeat the setup of previous § 4.1 with only human-annotated 595 relevant (query, document) pairs.
-We find the utilization of expert-informed definitions produces similar results, at most improving nDCG (see App.
-K).
-We argue that the definition might only make a significant difference when predicting all (query, document) pairs instead of only the human-annotated ones.
-The inclusion of specific nuances might especially help when compared with non-relevant documents.
-Calculating the nDCG and MAP score for all 43k (query, document) pairs, we find evidence for this notion (see Table 4).
-Thus, the inclusion of expert-informed definitions seems to improve the performance (for more details, see App.
-K).
-4.3 RQ3: Mitigating Annotation Bias
+RQ3: Mitigating Annotation Selection Bias.
 ClimRetrieve employs a real-world analyst scenario.
@@
 Therefore, the dataset allows us to investigate our model’s capabilities to counteract biases.
-For this purpose, we reannotated 288 strategically sampled (sampling details in App.
-M) (document, query) pairs to investigate Llama3-Tok’s performance on ClimRetrieve’s missing annotation.
-For samples not labeled as relevant in ClimRetrieve but by Llama3-Tok, the reannotations indicate that we can effectively overrule these blind spots.
-As Fig.
-4 shows, increasing the threshold of Llama3-Tok confidence allows us to be increasingly sure that we indeed find a relevant document.
-For the around 60% of the samples with a Llama3-Tok confidence higher than 99%, almost all samples are reannotated as relevant.
-Besides annotation selection bias, there are also other reasons for overruling the initial annotations.
-ClimRetrieve annotators may rather have focused on hard relevant documents omitting partially relevant ones.
-Furthermore, our GPT-4 created definitions may broaden the scope of what is relevant.
-However, we find that our model is well-calibrated and consistent with the provided relevance definitions.
-Therefore, Llama3-Tok is helpful for mitigating annotation selection bias (for a detailed analysis, see App.
-M).
-Figure 4: This table shows the percentage of agreement between Llama3-Tok and our human reannotation, and the amount of data remaining, when adjusting the confidence threshold for (query, document) pairs that were considered relevant by Llama3-Tok but not by the ClimRetrieve annotation.
-4.4 RQ4: Using $\mathcal{M}_{o}$ to Benchmark IR
+For this purpose, we sample 200 disagreements between Llama3-Tok’s annotation and the original ClimRetrieve labels.
+Then, we reannotate these samples with a human labler.
+To account for different confidence levels in Llama3-Tok’s prediction, we differentiate prediction with confidence higher or lower than 0.95.
+As Table 5 indicates, the model can be successfully used to overturn decisions of unseen, as irrelevant assumed documents (91.30% for confidence $>0.95$).
+Strikingly, even samples annotated by humans, i.e., those labeled as relevant, can be overturned, though with a lower certainty.
+We attribute this to differences in the unknown mental model of the ClimRetrieve labeler and our explicit relevance definitions (for details, see App.
+L).
+However, for us, it is reaffirming to observe that the DIRAS model is consistent with its own definition.
+Thus, we conclude that DIRAS’ labeling is effective and helps to mitigate annotation selection bias.
+ClimRetrieve (62.00%)
+All Rel. Irr.
+Conf $\leq 0.95$ 36.84 27.90 48.48
+Conf $>0.95$ 85.48 78.18 91.30
+Table 5: Accuracy of student model Llama3-Tok annotations that disagree with original ClimRetrieve relevance labels.
+All denotes all sampled disagreed labels.
+Rel.
+( Irr.) denotes the subset where the original label is relevant (irrelevant).
+Conf $\leq 0.95$ ($>0.95$) denotes the subset where Llama3-Tok’s confidence is lower (higher) than 0.95.
+ClimRetrieve (62.12%) means that 62.12% of data are annotated with $>0.95$ confidence.
 Setting Kendall’s $\mathbf{\tau}$
@@
 BGE-Large-ft 36.55
-Table 5: Different embedding models’ performance benchmarked by $\mathcal{M}_{o}$’s prediction on all 43K (query, document) pairs of ClimRetrieve.
+Table 6: Different embedding models’ performance benchmarked by student model $\mathcal{M}_{s}$’s prediction on all 43K (query, document) pairs of ClimRetrieve.
 “ft” denotes the model is fine-tuned on in-domain data.
-After validating $\mathcal{M}_{o}$’s performance, we can use it to annotate all 43k ClimRetrieve datapoints and obtain a benchmarking dataset to select IR algorithms.
+ELI5 ASQA QAMPARI RAG-Bench
+N N@10 N@20 N N@10 N@20 N N@10 N@20 N N@10 N@20
+GPT-4 48.43 15.97 17.81 64.62 38.82 46.32 56.21 27.54 35.34 42.91 31.84 41.35
+Llama3-Tok 48.73 17.04 20.08 64.90 39.37 48.44 56.58 28.70 35.49 48.27 41.13 47.88
+Table 7: Applying DIRAS to QA datasets, the IR performance of student model Llama3-Tok and GPT-4. N denotes nDCG.
+ELI5 (85.05%) ASQA (66.32%) QAMPARI (72.66%) RAG-Bench (60.33%)
+All Rel. Irr. All Rel. Irr. All Rel. Irr. All Rel. Irr.
+Conf $\leq 0.95$ 62.35 48.48 71.15 67.27 71.43 66.67 74.11 73.33 74.23 58.41 38.10 63.04
+Conf $>0.95$ 84.71 83.12 100.0 90.09 95.83 88.51 90.00 81.25 91.49 96.36 95.24 96.63
+Table 8: Accuracy of student model Llama3-Tok annotations that disagree with original relevance labels. Same setup as Table 5.
+RQ4: Benchmarking IR.
+We use $\mathcal{M}_{s}$ to annotate all 43K ClimRetrieve datapoints and obtain a benchmarking dataset to select IR algorithms.
 This approach can be especially helpful when lacking human annotation and annotation selection bias is prevalent.
 Specifically, we compare the performance of embedding models before and after in-domain fine-tuning.
-If the $\mathcal{M}_{o}$-annotated benchmark gives higher scores to the fine-tuned checkpoints, that means it is capable of selecting a better model for this specific domain.
-For this experiment, we first fine-tune open-sourced embedding models bge-large-en-v1.5 and bge-base-en-v1.5 (Chen et al., 2024) on ChatReport test set [^7] We fine-tune on the test instead of the training set to (1) leverage high-quality human annotation for fine-tuning; and (2) avoid indirect data leakage as $\mathcal{M}_{o}$ is fine-tuned on the training set.
+If the $\mathcal{M}_{s}$-annotated benchmark gives higher scores to the fine-tuned checkpoints, that means it is capable of selecting a better model for this specific domain.
+For this experiment, we first fine-tune open-sourced embedding models bge-large-en-v1.5 and bge-base-en-v1.5 (Chen et al., 2024) on ChatReportRetrieve test set [^6] We fine-tune on the test instead of the training set to (1) leverage high-quality human annotation for fine-tuning; and (2) avoid indirect data leakage as $\mathcal{M}_{s}$ is fine-tuned on the training set.
 (fine-tuning details in App.
-L).
+K).
 We then compare embedding models’ relevance ranking with the predicted ranking of Llama3-Tok on all 43K (query, document) pairs in ClimRetrieve.
@@
 This indicates the necessity of domain-specific benchmarking to tell the in-domain performance.
+4.2 Applying to QA Datasets
+In this section, we apply the DIRAS pipeline to QA datasets that are widely used in RAG benchmarking.
+DIRAS addresses queries for broad information and IR recall.
+Thus, we include long-form QA datasets from ALCE (Gao et al., 2023), including ELI5 (Fan et al., 2019), ASQA (Stelmakh et al., 2023), and QAMPARI (Amouyal et al., 2023).
+We also include RAG-Bench (Fang et al., 2024) that consists of questions from TriviaQA (Joshi et al., 2017), WebQ (Berant et al., 2013), and Natural Questions (Kwiatkowski et al., 2019).
+RAG-Bench is chosen since it has labels for partial vs.
+full relevance, which is a focus of DIRAS.
+Importantly, the context relevance labels from ALCE and RAG-Bench are derived from reference answers to questions, using heuristics instead of manual annotation.
+Thus, they are to some extent noisy.
+We first run the DIRAS pipeline [^7] Different from Climate change datasets, QA datasets do not require nuanced relevance definition.
+Thus we use a fixed relevance definition: “the document is helpful only if its content answers the query”.
+See full prompt in Fig.
+16.
+on each dataset to obtain corresponding Llama3-Tok models.
+Then we compare them with the teacher model – GPT-4.
+Results in Table 7 show that Llama3-Tok outperforms GPT-4 in IR.
+Then for each dataset, we repeat the annotation selection bias assessment of RQ3 in § 4.1.
+Again, we sample 200 disagreements between Llama3-Tok annotation and the original (potentially noisy) relevance labels, and manually check whether Llama3 or original labels are correct.
+As shown in Table 8, Llama3-Tok’s annotations are predominately correct with a confidence $>0.95$ (e.g., $85.05\%$ of ELI5).
+When there is a disagreement, relying on Llama3-Tok leads to less error (Acc.
+$>50\%$), especially when the confidence $>0.95$, thanks to the good calibration.
+For ASQA, QAMPARI, and RAG-Bench, the majority ($>90\%$) of the disagreement lies in originally irrelevant labeled part of the dataset (Irr.), possibly due to (query, document) pairs are selectively annotated.
+DIRAS achieves high accuracy in Irr.
+disagreements.
+Therefore, we reaffirm the notion that applying DIRAS to annotate broader (query, document) pairs can effectively reduce annotation selection bias, and thus improve IR recall benchmarking.
+All implementation details are in App.
+M.
 5 Recommendation for Future RAG
@@
 Therefore, we suggest not using top-k IR, avoiding the prior determined k that does not fit the actual amount of relevant information.
-Given the calibrated prediction of DIRAS fine-tuned $\mathcal{M}_{o}$, an alternative way is to retrieve all documents whose relevance scores exceed a pre-defined threshold.
+Given the calibrated prediction of DIRAS $\mathcal{M}_{s}$, an alternative way is to retrieve all documents whose relevance scores exceed a pre-defined threshold.
 Thus, different questions can retrieve different amounts of information depending on whether passing the relevance threshold.
@@
 This figure shows the change of F1 scores for obtaining relevant documents by thresholds.
-6 Related Literature
+6 Background and Related Work
 IR plays an important role in RAG but also becomes a performance bottleneck (Gao et al., 2024).
@@
 Low recall may leave out critical information for analysis (Ni et al., 2023).
-Prior work proposes various datasets and algorithms to benchmark RAG performance (Saad-Falcon et al., 2023; Chen et al., 2023a; Niu et al., 2024), but most of them focus on the final generation quality.
-No previous work studies how to efficiently generate domain-specific IR benchmarks for RAG.
-Sun et al.
+Domain-specific knowledge is also important for retrieval performance (Tang and Yang, 2024).
+Prior work on IR in RAG has already explored the idea of using LLMs to judge relevance.
+The closest to our work are RAGAs and ARES.
+Es et al.
+(2024) develop RAGAs to evaluate the relevance of each sentence in a paragraph with a closed-source LLM and create an aggregated score by dividing the number of relevant sentences over all sentences.
+Saad-Falcon et al.
+(2023b) aim to bring a relevance judge to a target domain through few-shot in-context learning (ICL): they first generate synthetic questions to given target-domain passages, and then fine-tune small classifiers for relevance judgment.
+However, many real-world challenges remain unaddressed.
+Specifically, IR recall, partial relevance, and domain-specific relevance definitions are neglected.
+ARES appears to have target-domain IR evaluation, but the synthetic data approach focuses on less integrative queries: each question is generated given a single passage and hard negatives are passages sampled from the same document, which could be (partially) relevant for integrative queries asking for broader information.
+Furthermore, we also find that few-shot ICL fails to teach domain-specific relevance to LLMs (§ 3.2).
+Besides, Sun et al.
 (2023b, a); Pradeep et al.
@@
 (2024) find that SOTA generic LLMs are good rerankers and such ability can be distilled to open-sourced LLMs.
-However, they all focus on pairwise or listwise ranking methods, while our work shows that the pointwise method (1) better fulfills the need of annotating domain-specific IR; and (2) works better with proper calibration method (Tian et al., 2023).
-Our work follows the stream of research annotating document relevance with LLMs (Thomas et al., 2024), taking one step further by (1) showing how to annotate with small LLMs; (2) predicting relevance scores to address partial relevance; and (3) use them to benefit RAG development.
+These studies all focus on pairwise or listwise ranking methods, and discusses that pointwise methods may not work due to bad calibration (Qin et al., 2024).
+However, when it comes to relevance annotation instead of reranking, pointwise relevance prediction is more suitable because it: (1) analyzes and annotates (query, document) pairs one-by-one, thus is more efficient and may better consider relevance definitions; and (2) can annotate both document rank and binary relevance labels (relevant or irrelevant) which are important for RAG in order to decide what information should be passed to the generator.
+We also show pointwise annotation works better with proper calibration method (Tian et al., 2023).
 7 Conclusion
 In this work, we introduce the DIRAS pipeline to fine-tune open-source LLMs to calibrated annotators.
-We prove the effectiveness of the approach on two dataset sets.
 The DIRAS approach has two significant advantages: (1) it is case-specialised allowing the incorporation of domain-specific knowledge into definitions, and (2) it helps to efficiently label a huge amount of documents with calibrated relevance scores.
@@
 As with every work, this has limitations.
-Our first limitation stems from the usage of two datasets focusing on a scenario of RAG report analyses.
-Given our expertise, this allows us to extend the deepness of our investigations: annotating domain-specific benchmarks and conducting error analyses.
-However, this limits the wideness of our research.
-While we argue that the results are representative of other knowledge-intensive RAG scenarios, it remains an open question for future work to generalize the DIRAS pipeline.
+First, our results show that DIRAS fine-tuning grants small student LLMs GPT-4-level performance on specific domains, but GPT-4 is not guaranteed to be perfect in all domains and cases.
+In certain niche domains, it might be necessary to augment GPT-4 with domain knowledge or agentic designs to achieve human-level performance in relevance annotation, and then create reliable training data for DIRAS.
+Although performance not always guaranteed, LLM annotation for document relevance is still necessary due to the sheer volume of (query, document) pairs and selection bias of human annotation.
 Second, this project focuses on text documents.
@@
 While this also presents a general limitation of modern-day RAG systems, we believe it is a crucial future step to generalize DIRAS’s idea of scalable information retrieval benchmarking to multi-modality.
-Our third limitation, and also a viable option to address multimodality, lies in the recent introduction of long-context LLMs.
+Our third limitation, and also a viable option to address multi-modality, lies in the recent introduction of long-context LLMs.
 These may make the role of information retrieval in RAG less crucial as entire documents can be used to answer a question.
@@
 Our integrated Quality, Environmental and Occupational Health & Safety (QEHS) Management System enables the achievement of harmonized K&S worldwide objectives.
-Furthermore, conclude as to whether this is relevant to answer the following question and definition:
+Furthermore, conclude as to whether this is relevant to answer the following query and definition:
 Meaning of the question: The question "What processes does the organization use to identify and assess climate-related risks?" is asking for information about the specific methods, tools, or strategies that a company employs to recognize and evaluate the potential risks to its operations, financial performance, and overall sustainability that are associated with climate change.
@@
 2. Engagement with external consultants or experts specializing in climate science [...]
-The question is clearly looking for processes to identify and assess risks associated with climate change.
+The query clearly looks for processes to identify and assess risks associated with climate change.
 Example 1.
@@
 The reason lies in the fact that when asking whether this document is relevant to the question, the answer is "partially right".
-Appendix B Creation of Question Relevance Definition
+Appendix B Creation of the Relevance Definition
 Fig.
-7 shows the prompt template for the creation of the question relevance definition.
+7 shows the prompt template for the creation of the query relevance definition.
 We ask the model to produce a short definition on which the model should rely.
@@
 For a demonstration of the output, see App.
-J.
-An analyst posts a <question> about a sustainability report.
-Your task is to explain the <question> in the context of sustainability reporting.
+I.
+An analyst posts a <question> about a climate report.
+Your task is to explain the <question> in the context of climate reporting.
 Please first explain the meaning of the <question>, i.e., the meaning of the question itself and the concepts mentioned.
-And then give a list of examples, showing what information from the sustainability report the analyst is looking for by posting this <question>.
+And then give a list of examples, showing what information from the climate report the analyst is looking for by posting this <question>.
 For <the questions meaning>, please start by repeating the question in the following format:
@@
 Examples of information that the question is looking for: <list of example information that the question is looking for>"""
-Figure 7: RAG Prompt Template enforcing structured output.
+Figure 7: Prompt for generating a query relevance definition.
 Appendix C Metrics Computation Details
@@
 For nDCG, relevant scores 0.5 and 1 are assigned to partially relevant and relevant documents correspondingly.
-Appendix D PDF Parsing and Document Length
-We use IBM deepsearch parser (Team, 2022) to parse corporate reports into chunks.
+When averaging Calibration metrics, we average AUROC with $1-$ ECE and $1-$ Brier Score to keep the trend consistent.
+Appendix D ChatReportRetrieve Data Preprocessing
+Data Sampling: We sample 31 questions and 80 reports from this application.
+Importantly, ChatReportRetrieve differs from general QA dataset and focuses on integrative queries which usually ask for broad relevant information.
+Therefore, with controlled annotation budget (i.e., number of (query, document) pairs), we are prone to have fewer (but representative) queries and more documents for each query.
+The queries are strategically sampled to ensure representativeness and diversity.
+Specifically, 11 queries are the core queries used in ChatReport, which cover essential topics of sustainability disclosure.
+20 questions are selected from users’ customized questions posed to the ChatReport tool.
+Climate reports are sampled randomly from openly accessible user submissions [^8] See https://github.com/EdisonNi-hku/chatreport.
+Finally, we prompt GPT-4 to draft relevance definitions for all queries (see App.
+B).
+PDF Parsing: We use IBM deepsearch parser (Team, 2022) to parse corporate reports into chunks.
 For chunks shorther than 120 tokens, we concatenate them with adjacent chunks to form chunks longer than 120.
@@
 8 shows the formatted chunks length distribution.
-Figure 8: Distribution of chunk length after being extracted from sustainability reports and concatenation.
+Figure 8: Distribution of chunk length after being extracted from climate reports and concatenation.
+Train-Test Split: We split the questions into 11 for testing and 20 for training.
+Similarly, we split the reports into 30 for testing and 50 for training.
+This ensures the evaluation on unseen queries and reports.
+For each query, we randomly sample 60 documents – 30 each from in the top-5 and outside the top-5 (using OpenAI text-embedding-3-small as the dense retriever).
+Ultimately, (query, document) pairs in training split are used to create training data with relevance label and confidence score predictions (details in § 3.2).
+Data points in the test split are passed to human annotation.
+We use 31 climate-related queries for data sampling among 80 climate reports.
+For the separation of the train dataset and test dataset, these queries are classified into 3 categories: vague, specific, and TCFD.
+Within each category, queries are split randomly for training and testing, resulting in 20 queries for the train set and 11 queries for the test set.
+Meanwhile, the 80 climate reports are randomly split into 50 for train and 30 for test.
 Appendix E Expert Annotation Process
-To obtain an expert-annotated test set, we annotate the test data that contains equally 330 samples of each top-5 relevant and non-to-5 relevant documents to the 11 questions/queries, thus obtaining 660 (query, document) pairs including the corresponding expert definitions.
 As described in App.
-D, the data is obtained from real sustainability reports and split into chunks of around 150 words with the IBM deepsearch parser (Team, 2022).
+D, the data is obtained from real climate reports and split into chunks of around 150 words with the IBM deepsearch parser (Team, 2022).
 Table 9 shows an overview of statistical properties of the number of words in test set data.
@@
 660 150 28.5 107 131 143 162 318
-Table 6: Statistical properties of the number of words in ChatReport test set data.
+Table 9: Statistical properties of the number of words in ChatReportRetrieve test set data.
 Then, we form a group of three expert annotators.
@@
 The definition consists of a question explanation and examples of relevant information.
-  • If a paragraph clearly fall into the definition of relevance, i.e., explicitly mentioned by the question explanation or examples, please annotate relevant.
+  • If a paragraph clearly falls into the definition of relevance, i.e., explicitly mentioned by the question explanation or examples, please annotate relevant.
   • If the paragraph is not explicitly covered by the definition but you think it somehow helps answering the question. Please annotate partially relevant.
   • Otherwise please annotate irrelevant.
-Additionally, one PhD student focusing on climate change and sustainability research serves as a subject-matter meta annotator to resolve conflicts or investigate cases where both labelers arise at the label partially.
+Additionally, one PhD student focusing on climate change and sustainability research serves as a subject-matter meta-annotator to resolve conflicts or investigate cases where both labelers arise at the label partially.
 Comparing the two base annotators in the setup, we can calculate inter-annotator agreement.
@@
 Not Relevant 474
-Table 7: Label distribution in the ChatReport test dataset.
+Table 10: Label distribution in the ChatReportRetrieve testset.
 Appendix F LLM Fine-Tuning Settings
@@
 Fig. 9 shows the full prompt DIRAS prompt template for the Chain-of-Thought setup. The non-CoT setup just excludes the “[Reason]: …"" part of the prompt.
-You are a helpful assistant who assists human analysts in identifying useful information within sustainability reports for their analysis.
+You are a helpful assistant who assists human analysts in identifying useful information within climate reports for their analysis.
 You will be provided with a <question> the analyst seeks to answer, a <paragraph> extracted from a lengthy report, and <background_information> that explains the <question>.
@@
 Figure 13: Listwise prompt with an extra input of explicit definition.
-Appendix I ClimRetrieve Dataset Overview
-The ClimRetrieve dataset simulates the typical tasks of a sustainability analyst.
-The annotators examine 30 sustainability reports.
-We use the report-level dataset of this paper which contains 43K (query, document) pairs labelled by relevance.
-This dataset is very long because every report is repeated per question to have unique (query, document) pairs.
-We also use the much shorter, only relevant (query, document) pairs containing 595 unique samples.
-Since the annotators only search for relevant information, these are the gold labels.
-There is no active labeling of the irrelevant (query, document) pairs.
-Table 8 offers a comparison of the statistical properties of the word count of the documents in ClimRetrieve.
-They are slightly longer than the ChatReport data (compare Table 6).
-Table 9 shows the label distribution in the relevant-only dataset.
-It becomes apparent that most of the (query, document) pairs are very relevant (label=3).
-This aligns with the report analyst setting where information is searched for until the question can effectively be answered.
-Number of words per document
-Dataset Dataset Size Mean Std Min 25% 50% 75% Max
-Report-Level 43.445 172 60.0 1 135 186 220 499
-Only Relevant 595 199 43.8 38 177 213 228 281
-Table 8: Statistical properties of the number of words in ClimRetrieve data.
-Label Occurance
-1 100
-2 167
-3 328
-Table 9: Distribution of relevance labels over the relevant (query, document) pairs in ClimRetrieve.
-Appendix J Creation of the Expert-Informed Relevance Definition
+Appendix I Creation of the Improved Relevance Definitions
 Fig.
-14 shows the prompt for the creation process of the expert-informed relevance definitions.
+14 shows the prompt for the creation process of the improved relevance definitions.
 Following the procedure in Schimanski et al.
@@
 Similar to the base setup for the experiments in Schimanski et al.
-(2024b), we use the text samples with a score of 2 or higher to create the expert-informed relevance definition.
-We include every relevant text part as an example.
-Thus, the prompt includes a large set of examples per question that can be synthesized in the relevance definition.
-While this procedure creates an expert-informed definition, it also represents a data leakage.
-In a sense, the definition will retrofit with the documents it searches for.
-While this is a limitation, we argue that this is also the step an expert human would take when annotating.
-She will have fixed concepts in her head, maybe even inspired by prior search processes.
-Therefore, we argue that this data leakage experiment is still adequate to investigate whether an expert-informed definition helps align the search process.
-Plugging the examples into the prompt results in a set of expert-informed relevance relevance definitions.
+(2024b), we use the text samples with a score of 2 or higher to create the improved relevance definition.
+We include relevant text samples in the prompt for creating the relevance definitions to obtain improved definitions.
+The logic behind this definition creation is that we assume we know the mental model of ClimRetrieve human analysts, and thus know what information is relevant before annotation.
+This is common in corporate report analysis where experts will have fixed concepts in their heads, maybe even inspired by prior search processes.
+Plugging the examples into the prompt results in a set of improved relevance definitions.
 When comparing these relevance definitions to the generic ones, it becomes apparent that GPT-4 already incorporated the majority of the concepts that the experts were looking for.
@@
 One example is displayed in App.
-J.
+I.
 While the meaning of the question remains rather static, there are nuanced differences in the examples that guide the relevance labeling.
-An analyst posts a <question> about a sustainability report.
-Your task is to explain the <question> in the context of sustainability reporting.
+An analyst posts a <question> about a climate report.
+Your task is to explain the <question> in the context of climate reporting.
 Please first explain the meaning of the <question>, i.e., meaning of the question itself and the concepts mentioned.
-And then give a list of examples, showing what information from the sustainability report the analyst is looking for by posting this <question>.
+And then give a list of examples, showing what information from the climate report the analyst is looking for by posting this <question>.
 For <the questions meaning>, please start by repeating the question in the following format:
@@
 Examples of information that the question is looking for: <list of example information that the question is looking for>"""
-Figure 14: RAG Prompt Template enforcing structured output with the inclusion of examples.
+Figure 14: Prompt Template enforcing structured output with the inclusion of examples.
 Question Generic Definition Expert-informed Definition
 Does the company provide definitions for climate change adaptation?
-"Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly outlined what it understands by the term ""climate change adaptation"" within its sustainability reporting.
+"Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly outlined what it understands by the term ""climate change adaptation"" within its climate reporting.
 This involves understanding how the company perceives the process of adjusting to actual or expected climate change and its effects, as well as how it plans to manage or mitigate harm or exploit beneficial opportunities.
@@
 6.
-Insights into how the company measures or evaluates the effectiveness of its adaptation strategies." "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly defined what it means by climate change adaptation within its sustainability reporting.
+Insights into how the company measures or evaluates the effectiveness of its adaptation strategies." "Meaning of the question: The question ""Does the company provide definitions for climate change adaptation?"" is asking for information about whether the company has clearly defined what it means by climate change adaptation within its climate reporting.
 This involves understanding how the company interprets climate change adaptation, including any specific strategies, initiatives, or policies it has in place to adjust to current or expected future climate change impacts.
@@
 Table 11: Example of a generic and expert-informed relevance definition for a question.
-Appendix K MAP and nDCG Scores for Different Relevance Definitions
-In the expert-informed definition experiment, we compare two settings.
+Appendix J MAP and nDCG Scores for Different Relevance Definitions
+In the improved definition experiment, we compare two settings.
 First, we compare the predictions on the 595 relevant-only (query, document) pairs.
-This is a replication of the setting in § 4.1.
-Since we don’t have non-relevant samples, we can only compare the nDCG.
+This is a replication of the setting in LABEL:sec:climre_veri.
+Since we do not have non-relevant samples, we can only compare the nDCG.
 Table 12 shows the results.
-It becomes apparent that only for the general nDCG score, the expert-informed query rates better.
+It becomes apparent that only for the general nDCG score, the improved query performs better.
 For the nDCG@5, and nDCG@10, the best-performing model remains with the generic prompt.
@@
 This could be a result of the definition creation.
-We use examples of relevance labels 2 and 3 to create the expert-informed definition.
+We use examples of relevance labels 2 and 3 to create the improved definition with GPT-4.
 Thus, we implicitly equalize relevance 2 and 3 in importance.
 This means we are likely less effective in differentiating between 2 and 3.
-This could explain the lower results for at lower k’s where differentiating between 2 and 3 is important v.s.
+This could explain the lower results at lower k’s where differentiating between 2 and 3 is important v.s.
 the overall nDCG where differentiating between 1 and 2/3 plays a more important role.
@@
 Table 13: Comparison of using the generic and the expert-informed relevance definitions for ranking all ClimRetrieve (query, document) pairs.
-Appendix L Embedding Fine-Tuning
+Appendix K Embedding Fine-Tuning
 We follow the official fine-tuning example [^10] https://github.com/FlagOpen/FlagEmbedding/tree/master/examples/finetune of (Chen et al., 2024) to fine-tune the embedding models.
-The models are fine-tuned on all annotated (query, document) pairs in ChatReport test set for 10 epochs, with a batch size of 4.
+The models are fine-tuned on all annotated (query, document) pairs in ChatReportRetrieve test set for 10 epochs, with a batch size of 4.
 Other hyperparameters are the same as the official example.
-Appendix M Hand-Checking of Model Results on ClimRetrieve
-The ClimRetrieve dataset presents the opportunity to investigate the output of our best-performing fine-tuned model Llama3 without CoT from different angles.
-In this experiment, we hand-annotate the output of our model in different categories and compare it to the ClimRetrieve annotations.
-As outlined in § 4, one issue of our approach is that we can only approximate the mental model of the annotator to create the relevance definitions.
-For the analysis, we consider the predictions of the fine-tuned Llama-3 model with the expert-informed relevance definitions created in App.
-J as they likely present the closest approximation to the mental model of the labelers.
-Furthermore, the ClimRetrieve dataset only has golden labels for the (partially) relevant documents.
-This has implications for our results.
-In this setup, we view our model as a calibrated annotator acting according to the given relevance definition – i.e., our model represents a fictional golden truth.
-This allows us to perform a qualitative edge case analysis on the ClimRetrieve data in different categories.
-The first category is the true positive classifications (model says relevant, ClimRetrieve-human says relevant).
-Since they are golden annotations by the ClimRetrieve annotators and align with our model, they are not checked as they likely are error-free.
-A more nuanced view has to be adopted for the three categories false negatives (model says relevant, ClimRetrieve-human says not relevant), true negatives (model says not relevant, human says not ClimRetrieve-relevant), and false positives (model says not relevant, human says ClimRetrieve-relevant).
-To obtain a qualitative understanding, we investigate the appearances (see Table 13) and confidences (see Table 14) of the categories.
-For the confidences, we use the empirically best-performing model in information retrieval Llama3-Tok (see Table 2).
-To perform a qualitative investigation, we sample (document, query) pairs from each category and reannotate them as being "relevant", "partially relevant" or "not relevant".
-These reannotations do not serve to obtain a holistic view (as already done in § 3) but rather gain insights into the special cases of the model’s predictions.
-First, we investigate the false negatives (model says relevant, human says not relevant).
-Investigating these cases is of particular interest since we know that humans have a selection bias (Thakur et al., 2021).
-In the analyst scenario setup of ClimRetrieve, this circumstance is aggravated by the fact that analysts may only search for information until they deem that they can answer the question at hand.
-Looking at Table 13, it becomes apparent that there is a very large number of false negatives.
-However, Table 14 shows that the confidence in the true negatives is much higher than in the false negatives.
-This indicates that the false negatives contain a much larger spectrum of partial relevance.
-This can be an initial explanation for the inflated number of false negatives.
-To perform a qualitative investigation, we sample (document, query) pairs from the false negatives and perform a reannotation with a single annotator.
-We have 16 unique questions/queries in the ClimRetrieve dataset and sample 6 (document, query) pairs per question/query.
-Furthermore, we want to investigate which role model confidence plays.
-Thus, the (document, query) pairs are sampled to contain two samples with a Llama3-Ask confidence above 0.9, two between 0.9 and 0.7, and two below 0.7.
-We choose Llama3-Ask confidence because it is easier to form thresholds and approximates the confidence well enough.
-Since we know that our model and the ClimRetrieve annotation inherently disagree, we focus on our reannotation.
-Fig.
-16 shows the resulting percentage of agreement between the model and our reannotation for increasing the confidence threshold.
-It becomes apparent that irrespective of the threshold, the agreement between our reannotator and the model is very high.
-Furthermore, increasing the threshold correlates with a high model-reannotation agreement.
-This indicates that the model’s assignment of the "relevant" label indeed works.
-At the same time, it bears the question of whether the initial ClimRetrieve-human annotation is invalid.
-Why does it underestimate the true relevant documents by such a large magnitude?
-There are three avenues to explain these results.
-First, as already mentioned, humans may miss out on sources, suffer from selection bias, not taking duplicate information into account, and may stop annotating once they reach a sufficient level of information for a question.
-While these things are inherent in the analyst setting of ClimRetrieve and out of our control, we can observe that ClimRetrieve clearly misses out on some relevant information (see for instance Fig.
-15).
-However and second, one further major reason for the inflated number of false negatives may be the change of definition we make by creating it with GPT-4.
-The original questions all concern the specific topic of climate change adaptation.
-This is a very narrow, specialized case of the general climate change domain.
-In our definitions, this generally looks different.
-Consider for instance the following example:
-Meaning of the question: The question "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?" is asking for information about whether the companys stated goals or objectives for environmental sustainability or climate change mitigation are aligned with, or make reference to, established external goals or targets.
+Appendix L Hand-Checking DIRAS Model Disagreements with ClimRetrieve
+We want to check samples where our student model Llama3-Tok’s prediction differs from the relevance label in ClimRetrieve.
+We sample equally from those samples where Llama3-Tok indicated relevance and ClimRetrieve does not and vice versa.
+While our focus lies on those documents that were not annotated in ClimRetrieve (i.e.
+irrelevant ones), the dataset also allows us to investigate how our model performs in the edge case of when a ClimRetrieve annotator deems relevance and Llama3-Tok does not.
+On the samples that were labeled as irrelevant in ClimRetrieve, we find that our student model Llama3-Tok model is effective in mitigating annotation selection bias and therefore incorporates the perspective of IR recall (see Table 5).
+However, it is interesting that our student model Llama3-Tok successfully overrules human decisions.
+We attribute this to the fact that our created relevance definitions might differ from the mental model of the human annotator in ClimeRetrieve.
+Thus, humans in ClimRetrieve might have been consistent with their own mental model.
+For us, however, it is more important and reaffirming to observe that Llama3-Tok is consistent with its own, explicit relevance definitions.
+We can view the (query, definition, document) pair in Fig.
+15 as an example.
+When analyzing the query "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?", the ClimRetrieve labeler interpreted the question broader, i.e., deeming this as relevant: "As a global technology leader, we are also committed to helping build the enabling societal conditions that will support a net zero economy.".
+However, for our student model, it is in line with the definition to assign a "not relevant" label.
+There is no explicit standard mentioned in the document.
+QUERY: "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?"
+QUERY DEFINITION: Meaning of the question: The question "Do the environmental/sustainability targets set by the company reference external climate change adaptation goals/targets?" is asking for information about whether the companys stated goals or objectives for environmental sustainability or climate change mitigation are aligned with, or make reference to, established external goals or targets.
 These external references could include international agreements, national policies, or standards set by recognized organizations focused on climate change and sustainability.
@@
 1. In line with our commitment to the Net-Zero Banking Alliance (NZBA) [...]
-While the question itself only addresses "climate change adaptation", the relevance definition allows for contents that are in line with "climate change mitigation".
-Mitigation is much broader and much more discussed in sustainability reports.
-The examples in the definition additionally broaden the scope.
-This also hints at the third avenue to explain the many false negatives: they contain a lot of partially relevant (document, query) pairs.
-This can also explain the significant drop in average confidence between true and false negatives in Table 14.
-However, the important aspect of the DIRAS pipeline is that it is consistent with the provided definitions which Fig.
-16 indicates.
-This investigation unveils clear edge cases for partial relevance and the mismatch of different definitions, i.e.
-mental models of annotators.
-Nonetheless, Fig.
-16 reveals that, even amongst these hard examples, the approach remains consistent with the provided definitions.
-Although looking at edge cases, our reannotations are largely aligned with our model when overturning the non-golden irrelevant documents of ClimRetrieve.
-To further get an intuition on edge cases, we turn our attention to the false positives.
-These represent the most complicated cases where the ClimRetrieve human annotator deemed the (document, query) pair relevant and the model arose with the judgment of not relevant.
-Since only 148 (document, query) pairs are in the false positives, we randomly sample 96 (document, query) pairs and reannotate them.
-The first two quantitative observations remain similar to the false negatives.
-First, Table 14 shows that the average confidence for the false positives is significantly lower than those of the true positives indicating they present edge cases.
-Indeed, they have the lowest average confidence among all categories with 0.8871.
-However, Fig.
-17 again shows that with a rising confidence threshold, the agreement between the prediction and our reannotation rises.
-This time, the level of partial relevance is higher than for the true negatives.
-These two cues towards higher complexity continue in the qualitative assessment of our reannotations.
-We repeat to ask the question why is there a discrepancy between our model and the ClimRetrieve annotations?
-This time, we identify a multitude of edge cases.
-First, a technical edge case comes into play.
-ClimeRetrieve was created by matching span-labeled relevant texts with documents retrieved from the reports.
-A document is deemed relevant when one sentence of the span-labeled text appears in a document.
-However, one sentence alone cut off or in a different context may not be relevant.
-In 19% (18 of 96) of the annotated samples, this is the reason for a wrong label assigned.
-Second, the nature of partially relevant labels seems to be ambiguous.
-While the ClimRetrieve human annotator deemed many samples partially relevant, our reannotation overturned some of these cases in line with our model.
-Nonetheless, there remains a significant chunk of partially relevant samples (see Fig.
-17).
-Third, using our model, we overturn some of the decisions of the ClimRetrieve annotators entirely.
-Even if they deemed it relevant, our reannotation and model suggest irrelevance given the definition.
-This can again be linked to the mismatch of the exact mental model of the ClimRetrieve annotators and our approximated reannotation guidelines.
-It has to be noted that these decisions may again be overruled by a second or third reannotator.
-The samples generally all have a connection with the topics in the question.
-Rather than presenting a gold standard, this comparison allows us to understand the fine-grained nuances of relevance, therefore reinforcing the very need of this project.
-Clear misannotations were only marked in 3 of 96 cases where our annotator could not identify any remote relevance.
-Finally, we investigate the large chunk of the (document, query) pairs, the true negatives (model says not relevant, human says not relevant).
-Having a large basis of 37.409 (document, query) pairs, we again sample 6 pairs per question using two samples with a Llama3-Ask confidence above 0.9, two between 0.9 and 0.7, and two below 0.7.
-The true negatives are not at all comparable with the investigations before because we now sample from pairs with apriori agreement.
-Thus, we don’t investigate edge cases in this category.
-This is also confirmed by the results (see Fig.
-18).
-Only two samples with a Llama3-Tok confidence threshold below 0.5 were labeled as partially relevant, none were labeled as relevant.
-This reinforces the general results in § 3 and shows that the previous two reannotations concerned edge cases.
-Collectively, these observations prove the motivation for this project in developing a nuanced stand beyond binary and even partial relevance.
-They show that the models function consistently with themselves - even in edge-case scenarios.
-This can mean something different than consistent with human annotators who cannot share their mental model.
-"Golden" model with
-"golden" definition
-True False
-Human with own mental model True
-447
-148
-False
-5.441
-37.409
-Table 13: Comparison of the Llama-3’s relevance prediction vs. the ClimRetrieve human annotators on ClimRetrieve’s 43K (document, query) pairs.
-Labels assigned by
-fine-tuned LLama3
-Average Confidence
-Llama3-Tok
-all 0.9681
-positives 0.9049
-negatives 0.978
-true positives 0.9575
-false positives 0.8871
-true negatives 0.9784
-false negatives 0.9006
-Table 14: Average confidence scores of Llama3-Tok for the different classification categories.
-relevance definition:
-Meaning of the question: The question "Has the company identified any synergies between its climate change adaptation goals and other business goals?" is asking for information about how the companys efforts to adapt to climate change are aligned with, or can complement, its other business objectives.
-This includes understanding if climate change initiatives also support broader strategic goals such as cost reduction, risk management, innovation, market expansion, or reputation enhancement.
-Examples of information that the question is looking for:
-1. Descriptions of [...]
-MISSED OUT DOCUMENT:
-[...] Along with these efforts, in each business segment, Sony develops and enhances risk management and business continuity plans (BCPs) from the perspective of improving risk management across supply chains, through the identification, analysis, and assessment of business continuity risks.
-Flood damage has grown in recent years due to the impact of climate change, prompting Sony to reassess the flood risk at its manufacturing sites in Japan and implement preventative measures that will mitigate flood damage and facilitate rapid recovery.
-Sony is collaborating with relevant companies and organizations, and conducts hands-on drills to address foreseeable risks, in an effort to enhance business continuity and accelerate flood recovery.
-Sony will continue to increase its resilience to climate change, based on its analyses and initiatives.* A global initiative in which participating corporations aim to operate on 100% renewable electricity.
-It is headed by an international non-governmental organization, the Climate Group, in partnership with the CDP.
-Figure 15: False Negative example of clearly relevant but missed out information in ClimRetrieve.
-Figure 16: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for false negatives.
-Figure 17: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for false positives.
-Figure 18: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for true negatives.
+DOCUMENT: Enabling a more sustainable world Microsofts actions alone will not solve the climate crisis.
+As a global technology leader, we are also committed to helping build the enabling societal conditions that will support a net zero economy.
+Were focused on accelerating the availability of new climate technologies, strengthening our climate policy agenda, helping to develop a more reliable and interoperable carbon accounting system, advocating for skilling programs to expand the green workforce, and working to enable a just energy transition.
+Figure 15: Example for which DIRAS Llama3-Tok assigns "not relevant" and ClimRetrieve assigned "relevant".
+This example shows that while DIRAS Llama3-Tok might differ with the opinion of ClimRetrieve’s annotator, it is consistent with its own definition.
+Appendix M Implementation Details of Experiments on QA Datasets
+ALCE Data: We obtain ELI5, ASQA, and QAMPARI from ALCE (Gao et al., 2023), where they parse the original open-domain QAs into RAG forms [^11] Data files in https://github.com/princeton-nlp/ALCE.
+For each question, ALCE annotates 5 documents as oracle based on retrieval recall and reference answers, which are used as context relevance labels in our experiment.
+For each dataset, we randomly sample 100 queries to construct DIRAS training data, following the process in Fig.
+2.
+Top-5 is selected for balanced sampling, thus resulting in 1000 (query, document) pairs for training.
+We sample 50 questions for test data, and include all (query, document) pairs for them, resulting in 5K (query, document) pairs for each dataset.
+RAG-Bench Data: RAG-Bench classifies RAG sources into four types: (A) relevant and with answers, (B) relevant topic but without answers, (C) irrelevant topic, and (D) with counterfactual answers.
+We find (B) addresses partial relevance that DIRAS cares about.
+Therefore, we leverage its dev set for training and test set for testing, where (A) and (D) become relevant documents, and (B) and (C) are used as irrelevant ones.
+Disagreement Sampling for Table 8: We sample 200 disagreed annotations four each dataset, 50 samples from each confidence range: Conf$<90$, $90<$Conf$<95$, $95<$Conf$<98$, and $98<$Conf$<100$ to balancedly cover different confidence scores (these bins are of similar size).
+You will be provided with a <question> and a <paragraph>.
+Your job is to assess whether the <paragraph> is useful in answering the <question>, given the <background_information> defining what is useful.
+<background_information>: "The <paragraph> is useful only if some of its content directly answer the <question> or at least a part of the <question>.
+Content with relevant topic but without direct answers are not useful."
+<question>: "{question}"
+<paragraph>: "{paragraph_chunk}"
+Is <paragraph> useful for answering <question>? Provide your best guess and your confidence that the guess is correct. Reply in the following format:
+[Reason]: <Reason why and how the paragraph is helpful or not helpful for answering the question. Clearly indicate your stance.>
+[Guess]: <Your most likely guess, should be one of "Yes" or "No".>
+[Confidence]: <Give your honest confidence score between 0.0 and 1.0 about the correctness of your guess.
+0 means your previous guess is very likely to be wrong, and 1 means you are very confident about the guess.>
+Figure 16: The DIRAS prompt for QA experiments.